### v3

# Multipoint Schwarz–Pick Lemma for the quaternionic case

Abstract.
Following ideas by Beardon, Minda and Baribeau, Rivard, Wegert in the context of the complex Schwarz–Pick Lemma, we use iterated hyperbolic difference quotients to prove a quaternionic multipoint Schwarz–Pick Lemma, in the context of the theory of slice regular functions. As applications, we obtain quaternionic Dieudonné and Goluzin estimates. Finally, an algorithm for the construction of (Nevanlinna–Pick) interpolating slice regular functions with real nodes is provided as a byproduct of the quaternionic multipoint Schwarz–Pick Lemma.
Key words and phrases:
Schwarz–Pick Lemma, slice regularity, Möbius transformations, hyperbolic difference quotient, Nevanlinna–Pick interpolation
2020 Mathematics Subject Classification:
30G35
The two authors were partially supported by GNSAGA of INdAM and by PRIN Varietà complesse: geometria, topologia e analisi armonica
Introduction
In complex analysis, holomorphicity plays an important role in the study of the intrinsic geometry of the unit disk $\mathbb{D}=\{z\in\mathbb{C}:|z|<1\}$, thanks to the Schwarz–Pick Lemma, which asserts the following:
Lemma 0.1 (Schwarz–Pick Lemma).
Let $f:\mathbb{D}\longrightarrow\mathbb{D}$ be a holomorphic function. Let $z_{0}\in\mathbb{D}$. Then for any $z\in\mathbb{D}$ it is
(1) $\left|\frac{f(z)-f(z_{0})}{1-\overline{f(z_{0})}f(z)}\right|\leq\left|\frac{z-z_{0}}{1-\overline{z_{0}}z}\right|;$
moreover
$\frac{|f^{\prime}(z_{0})|}{1-|f(z_{0})|^{2}}\leq\frac{1}{1-|z_{0}|^{2}}.$
Inequalities are strict for $z\neq z_{0}$, unless $f$ is an automorphism of the disk.
This lemma was formulated at the beginning of the 20th century by Pick [20] and it has also a more geometric interpretation: a holomorphic self-map of $\mathbb{D}$, if it is not an automorphism, is a contraction with respect to the Poincaré metric on $\mathbb{D}$, $ds^{2}=(1-|z^{2}|)^{-2}dzd\overline{z}$ and in the automorphism case it is an isometry. This result turns out to be a special case of Nevanlinna–Pick interpolation Theorem, which gives a necessary and sufficient condition for the existence of a holomorphic function $f:\mathbb{D}\longrightarrow\mathbb{D}$ such that $f(z_{m})=w_{m}$ for all $m=1,\ldots,n$, given $n$ distinct points $z_{1},\ldots,z_{n}\in\mathbb{D}$ and $w_{1},\ldots w_{n}\in\mathbb{D}$. The condition is that the Pick matrix
$P=\left(\frac{1-w_{a}\overline{w_{b}}}{1-z_{a}\overline{z_{b}}}\right)_{a,b=1,\ldots,n}$
is positive semidefinite. It is straightforward to see that when $n=2$ this condition reduces to inequality (1).
The Schwarz–Pick Lemma has been extended in various ways, in many years and by many authors. Among the others, Beardon and Minda [7] have introduced the following hyperbolic difference quotient
$f^{\star}_{z_{0}}(z):=\begin{cases}\frac{[f(z),f(z_{0})]}{[z,z_{0}]}&z\neq z_{0}\\ f^{\prime}(z_{0})\frac{1-|z_{0}|^{2}}{1-|f(z_{0})|^{2}}&z=z_{0},\end{cases}\qquad\text{where }[z,w]=\frac{z-w}{1-\overline{w}z}$
and with this quantity they have got a ‘three points’ version of the Schwarz–Pick Lemma
Theorem 0.2 ([7, Thm. 3.1]).
Let $f:\mathbb{D}\longrightarrow\mathbb{D}$ be holomorphic but not an automorphism. Let $z_{0},z_{1},z_{2}\in\mathbb{D}$. Then
$\left|\frac{f^{\star}_{z_{0}}(z_{1})-f^{\star}_{z_{0}}(z_{2})}{1-\overline{f^{\star}_{z_{0}}(z_{2})}f^{\star}_{z_{0}}(z_{1})}\right|\leq\left|\frac{z_{1}-z_{2}}{1-\overline{z_{2}}z_{1}}\right|.$
The inequality is strict for $z_{1}\neq z_{2}$ unless $f$ is a complex Blaschke product of degree two.
From this result they have also obtained easy proofs of classical estimates for derivatives of holomorphic self-maps of the unit disk, assuming that the origin is a fixed point.
Baribeau, Rivard and Wegert [6] have considered iterated hyperbolic difference quotients $f^{\{n\}}_{z_{1},\ldots,z_{n}}=(f^{\{n-1\}}_{z_{1},\ldots,z_{n-1}})^{\star}_{z_{n}}=(((f^{\star}_{z_{1}})^{\star}_{z_{2}})^{\star}_{z_{3}}\ldots)^{\ast}_{z_{n}}$, getting an analogous theorem with more than three points. Moreover, they used these iterated differences to give simpler conditions for the $n$ points Nevanlinna–Pick interpolation problem, and also a constructive algorithm for the solutions which is based on the works of Schur [21] and Nevanlinna [19]. Iterated hyperbolic difference quotients have been also considered in [13] in order to obtain estimates for higher order derivatives and recently in [1] to generalize Julia’s Lemma.
Over the skew-field of quaternions $\mathbb{H}$, which contains $\mathbb{C}$, the notion of slice regularity takes the place of holomorphicity. Slice regular functions have been introduced by Gentili and Struppa [17], developing a regularity concept that was originally formulated by Cullen [14]. It is based on fact that it can be given a ‘book structure’ to the skew-field of quaternions $\mathbb{H}=\bigcup_{I\in\mathbb{S}}\mathbb{C}_{I}$, where $\mathbb{S}$ is the sphere of imaginary units $\{q\in\mathbb{H}:q^{2}=-1\}$ and $\mathbb{C}_{I}$ are the complex subspaces $\{x+yI\in\mathbb{H}:x,y\in\mathbb{R}\}$, which are known as complex slices and which can be identified with complex planes having the real line $\mathbb{R}$ in common. A function $f:\Omega\longrightarrow\mathbb{H}$ defined on a domain in $\mathbb{H}$ is said to be slice regular if and only if its restriction to any slice $\mathbb{C}_{I}$ is holomorphic, i.e., the differential operator ¯$\overline{\partial}_{I}=1/2(\partial_{x}+I\partial_{y})$ vanishes on $f\mid_{\Omega\cap\mathbb{C}_{I}}$. If so, one can define a derivative $\partial_{C}f$ (the regular, or Cullen, derivative) in this way: $\partial_{C}f(q)=\partial_{I}(f\mid_{\Omega\cap\mathbb{C}_{I}})(q)$ for any $q\in\Omega\cap\mathbb{C}_{I}$, where $\partial_{I}$ is the operator $1/2(\partial_{x}-I\partial_{y})$. It is important to observe that the common pointwise product of functions does not preserve slice regularity; however, one could define a different product on spaces of slice regular functions, the $\ast$-product $f\ast g$, which allows to define on them the structure of a (noncommutative) algebra. Using this operation together with the $\ast$-inversion $f\mapsto f^{-\ast}$, one can define slice regular Möbius transformations on the quaternionic unit ball $\mathbb{B}=\{q:\mathbb{H}:|q|<1\}$, that is to say the maps of the form
$(1-q\overline{p})^{-\ast}\ast(q-p)u,\quad p\in\mathbb{B},u\in\partial\mathbb{B}.$
These above are the unique slice regular bijective functions mapping $\mathbb{B}$ onto itself. When $u=1$, we denote the map above by $\mathcal{M}_{p}$. Analogously, one can give a definition of slice regular finite Blaschke products on the unit ball $\mathbb{B}$, which can be characterized in a similar way as in the complex framework.
All the facts we need from the theory of slice regularity are reviewed in Section 1, with a focus on slice regular Möbius transformations (following [11, 22]) and regular Blaschke products on $\mathbb{B}$. For further information we refer to the recent monograph by Gentili, Stoppato and Struppa [16].
Bisi and Stoppato [10] gave a quaternionic version of the Schwarz–Pick Lemma in the framework of this theory. Analogously to what happens in the complex case, in Section 2 we extend the notion of hyperbolic difference quotient $f^{\star}_{p}$ for slice regular functions from $\mathbb{B}$ to itself, thus the quaternionic Schwarz–Pick Lemma leads to a three-points Schwarz–Pick Lemma:
Theorem 0.3.
Let $\mathbb{B}=\{q\in\mathbb{H}:|q|<1\}$ be the open quaternionic unit ball. Suppose $f:\mathbb{B}\longrightarrow\mathbb{B}$ is a slice regular function which is not bijective. Let $p,s\in\mathbb{B}$. Then for all $q\in\mathbb{B}$
$\left|(\mathcal{M}_{f^{\star}_{p}(s)}\bullet f^{\star}_{p})(q)\right|\leq\left|\mathcal{M}_{s}(q)\right|,$
where $\mathcal{M}_{f^{\star}_{p}(s)}\bullet f^{\star}_{p}$ stands for $(1-f^{\star}_{p}\ast\overline{f^{\star}_{p}(s)})^{-\ast}\ast(f^{\star}_{p}-f^{\star}_{p}(s))$. Equality holds for some $q\in\mathbb{B}\setminus\{s\}$ (and hence for any $q\in\mathbb{B}$) if and only if $f$ is a slice regular Blaschke product of degree two.
In Section 3 we show some consequences of this results. In particular, following [7] and restricting to the case of a fixed point at the origin, by the connection between the regular derivative $\partial_{C}f(p)$ and the hyperbolic derivative $f^{h}(p):=f^{\star}_{p}(p)$ for slice regular self-maps on $\mathbb{B}$, we get the quaternionic version of some inequalities well established in classical complex analysis, such as Dieudonné’s or Goluzin’s estimates (assuming $f(0)=0$).
The definition of iterated hyperbolic difference quotients in the quaternionic setting and the consequent multipoint Schwarz–Pick Lemma are presented in Section 4.
In Section 5 we consider the Nevanlinna–Pick interpolation problem with $n$ points in the quaternionic unit ball $\mathbb{B}$, looking for a slice regular interpolating function $f:\mathbb{B}\longrightarrow\mathbb{B}$ for the prescribed nodes and values. Here we give a generalization to the quaternionic setting of the conditions and the algorithm presented in [6] for the complex case, in terms of hyperbolic difference quotients. Since the quaternionic setting is wilder than the complex one and not commutative, we have to restrict the problem to the case of real nodes. Given $n$ distinct real points $r_{1},\ldots,r_{n}\in(-1,1)$ and $n$ values $s_{1},\ldots,s_{n}$ in the quaternionic unit ball $\mathbb{B}$, one defines by iteration quantities $Q_{\kappa}^{\ell}$ for $\kappa=1,\ldots,n-1$ and $\ell=\kappa+1,\ldots,n$ which are generically expressed as quotients of Möbius maps, starting from the given nodes and values. Then we have a criterion in terms of the last term $Q_{n-1}^{n}$:
Theorem 0.4.
There exists a slice regular function $f:\mathbb{B}\longrightarrow\mathbb{B}$ such that $f(r_{m})=s_{m}$ for all $m=1,\ldots,n$ if and only $Q_{n-1}^{n}$ is a quaternion of modulus less or equal to one. If $|Q_{n-1}^{n}|<1$, then there is an infinite family of solutions, while if $|Q_{n-1}^{n}|=1$ there exists a unique solution given by a regular Blaschke product.
The terms $Q_{\kappa}^{\ell}$ correspond to values of the iterated difference quotients of any solution of the problem: $Q_{\kappa}^{\ell}=f^{\{\kappa\}}_{r_{1},\ldots,r_{\kappa}}(r_{\ell})$. The algorithm used for proving this result also provides explicit formulas for the solutions.
When the nodes are not all real, the algorithm does not work anymore: we give some evidences in the last part of Section 5.
A quaternionic general Nevanlinna–Pick Theorem in terms of the positive semi-definition of the associated Pick matrix has already been proved by Alpay, Bolotnikov, Colombo and Sabadini in [2] with a different approach based on linear algebra and functional analysis techniques. Although this result is more general, it gives less explicit formulas for the solutions.
1. Preliminaries
1.1. Slice regular functions
Let $\mathbb{H}=\mathbb{R}+\mathbb{R}i+\mathbb{R}j+\mathbb{R}k$ denote the skew-field of quaternions. Let $\mathbb{S}$ be the sphere of imaginary units: $\mathbb{S}=\{q\in\mathbb{H}:q^{2}=-1\}=\{q\in\mathbb{H}:\operatorname{Re}(q)=0,|q|=1\}$. For any $I\in\mathbb{S}$, let $\mathbb{C}_{I}$ denote the complex subspace $\{x+yI\in\mathbb{H}:x,y\in\mathbb{R}\}$, which we will call a complex slice. We observe that $\mathbb{C}_{-I}=\mathbb{C}_{I}$ and that $\mathbb{H}$ is the union of such slices:
$\mathbb{H}=\bigcup_{I\in\mathbb{S}}\mathbb{C}_{I}.$
Moreover, all the slices have the real line in common: $\bigcap_{I\in\mathbb{S}}\mathbb{C}_{I}=\mathbb{R}$. This decomposition of the quaternionic space is sometimes known as book structure.
Definition 1.1.
Let $\Omega\subseteq\mathbb{H}$ be a domain. For any $I\in\mathbb{S}$, let $\Omega_{I}:=\Omega\cap\mathbb{C}_{I}=\{x+yI\in\Omega:x,y\in\mathbb{R}\}$. A function $f:\Omega\longrightarrow\mathbb{H}$ is said to be (left) slice regular on $\Omega$ ($f\in\mathcal{SR}(\Omega)$) if it has continuous partial derivatives and any restriction $f_{I}=f\mid_{\Omega_{I}}$ for $I\in\mathbb{S}$ is holomorphic, in the sense that ¯$\overline{\partial}_{I}f_{I}(x+yI)\equiv 0$ on $\Omega_{I}$, where ¯$\overline{\partial}_{I}$ is the differential operator
¯$\overline{\partial}_{I}:=\phi_{I}\mapsto\frac{1}{2}\left(\partial_{x}\phi_{I}+I\partial_{y}\phi_{I}\right),\qquad\phi_{I}:\Omega_{I}\longrightarrow\mathbb{H}.$
A trivial consequence of this definition, observing that $\mathbb{H}=\mathbb{C}_{I}+\mathbb{C}_{I}J$ for any $I\in\mathbb{S}$ and $J\in\mathbb{S}$ orthogonal to $I$ with respect to the standard scalar product of $\mathbb{R}^{4}$, is the following:
Lemma 1.2 (Splitting Lemma, [16, Lemma 1.3]).
Let $f:\Omega\longrightarrow\mathbb{H}$ be slice regular on a domain $\Omega\subseteq\mathbb{H}$. For any $I\in\mathbb{S}$ and $J\in\mathbb{S}$ such that $J\perp I$, there exist two holomorphic functions $F,G:\Omega_{I}\longrightarrow\mathbb{H}$, depending on $I$ and $J$, such that the restriction $f_{I}=f\mid_{\Omega_{I}}$ writes down as
$f_{I}(z)=F(z)+G(z)J\qquad\text{for all }z=x+yI\in\Omega_{I}.$
Definition 1.3.
Let $\Omega\subset\mathbb{H}$ be a domain and let $f:\Omega\longrightarrow\mathbb{H}$ be a slice regular function. For any $I\in\mathbb{S}$ we set $\partial_{I}f_{I}(x+yI):=\frac{1}{2}\left[\partial_{x}f_{I}(x+yI)-I\partial_{y}f_{I}(x+yI)\right]$. The regular derivative (or Cullen derivative) of $f$ is defined as
$\partial_{C}f(q):=\partial_{I}f_{I}(q)\quad\text{ if }q\in\Omega_{I}.$
The definition is well-posed since if $x_{0}\in\bigcap_{I\in\mathbb{S}}\Omega_{I}=\Omega\cap\mathbb{R}$, then $\partial_{I}f_{I}(x_{0})=\partial_{x}f(x_{0})$ for any imaginary unit $I$.
It is clear by the definition that the Cullen derivative of a slice regular function $f$ is again slice regular on the same domain, hence one can define higher (Cullen) derivatives $f^{(n)}(q)$ as well:
$f^{(n)}(q):=(\partial_{I})^{n}f_{I}(q)\quad\text{ if }q\in\Omega_{I}.$
Example 1.4.
It is easy to check that all polynomial functions with right quaternionic coefficient
$f(q)=q^{N}a_{N}+q^{N-1}a_{N-1}+\ldots+qa_{1}+a_{0}$
are slice regular functions over $\mathbb{H}$; moreover the power series of the form
$g(q)=\sum_{m=0}^{+\infty}q^{m}a_{m}$
are slice regular function over the ball of convergence $B(0,\varrho):=\{q\in\mathbb{H}:|q|<\varrho\}$, where $\varrho^{-1}=\limsup\limits_{m\to+\infty}|a_{m}|^{1/m}$.
On the other side, the function $q\mapsto i\cdot q$ is not slice regular, since if we consider the slice $\mathbb{C}_{j}$, then the relative restriction is $x+yj\mapsto xi+yk$ and ¯$\overline{\partial}_{j}(xi+yk)=\frac{1}{2}(i+jk)=\frac{2i}{2}=i$.
Remark 1.5.
The sum $f+g$ of two slice regular functions $f,g\in\mathcal{SR}(\Omega)$ is still slice regular. The last example shows that the standard pointwise product $fg$ of two slice regular functions $f,g$ is not always slice regular. Moreover, the composition of two slice regular functions $g\circ f$, if defined, is not in general slice regular: for instance $q\mapsto(q\cdot i)^{2}=q\cdot i\cdot q\cdot i$ is such that its restriction to $\mathbb{C}_{j}$ is $(x+yj)\mapsto(xi-yk)^{2}=-x^{2}-y^{2}$ and ¯$\overline{\partial}_{j}(-x^{2}-y^{2})=-(x+yj)\neq 0$ for all $x+yj\neq 0$.
Definition 1.6.
A domain $\Omega\subset\mathbb{H}$ is said a slice domain if $\Omega\cap\mathbb{R}\neq\varnothing$ and any intersection $\Omega_{I}$ with a slice $\mathbb{C}_{I}$ is connected.
A domain $\Omega$ as above is axially symmetric if for any of its points $x+yI$ (where $x,y\in\mathbb{R}$, $I\in\mathbb{S}$) the whole sphere $x+y\mathbb{S}:=\{x+yJ:J\in\mathbb{S}\}$ is contained in $\Omega$.
Making both these assumptions on the domain, we have the following remarkable representation formula for slice regular function.
Proposition 1.7 (Representation Formula, [16, Thm. 1.16]).
Let $f:\Omega\longrightarrow\mathbb{H}$ be a slice regular function defined on an axially symmetric slice domain Let $I,J,K\in\mathbb{S}$ such that $J\neq K$. Then
$f(x+yI)=(J-K)^{-1}[Jf(x+yJ)-Kf(x+yK)]+I(J-K)^{-1}[f(x+yJ)-f(x+yK)].$
In particular, if we choose $K=-J$, we get
(2) $f(x+yI)=\frac{1}{2}[f(x+yJ)+f(x-yJ)]+\frac{IJ}{2}[f(x-yJ)-f(x+yJ)].$
It is important to observe that formula (2) says that the values of $f$ on a slice determine the whole function.
If we consider balls centered in the origin as domains, we get that all functions which are slice regular can be written as power series:
Theorem 1.8 ([16, Thm. 1.10]).
Let $\Omega=B(0,R):=\{q\in\mathbb{H}:|q|<R\}$ for some $R>0$. Then for any $f:\Omega\longrightarrow\mathbb{H}$ slice regular we have
$f(q)=\sum_{m=0}^{+\infty}q^{m}(m!)^{-1}f^{(m)}(0).$
It can be defined a product between such functions which preserves slice regularity:
Definition 1.9.
Let $f(q)=\sum_{m=0}^{+\infty}q^{m}a_{m}$ and $g(q)=\sum_{m=0}^{+\infty}q^{m}b_{m}$ be two slice regular functions on a ball $B=B(0,R)$. Their $\ast$-product $f\ast g$ is the slice regular function on $B$ given by
$(f\ast g)(q):=\sum_{m=0}^{+\infty}q^{m}\left(\sum_{\ell=0}^{m}a_{m-\ell}b_{\ell}\right).$
This power series still converges in $B(0,R)$. Since the product in $\mathbb{H}$ is associative, then also the $\ast$-product is associative; however, it is clearly noncommutative.
In order to evaluate at a single point the $\ast$-product, we have the following formula.
Proposition 1.10 ([16, p. 42]).
For $f,g$ as in the previous definition, it is
$(f\ast g)(q)=\begin{cases}0&\text{if }f(q)=0\\ f(q)g(f(q)^{-1}qf(q)),&\text{otherwise.}\end{cases}$
The $\ast$-product, together with the standard sum, gives to $\mathcal{SR}(B(0,R))$ the structure of a (noncommutative) $\mathbb{R}$-algebra.
Definition 1.11.
A function $f:\Omega\longrightarrow\mathbb{H}$ is slice preserving if for any $I\in\mathbb{S}$ it is $f(\Omega_{I})\subseteq\mathbb{C}_{I}$.
It is one slice preserving if there exists a unique $I\in\mathbb{S}$ so that $f$ is $\mathbb{C}_{I}$-preserving: $f(\Omega_{I})\subseteq\mathbb{C}_{I}$.
Proposition 1.12.
Let $f\in\mathcal{SR}(B(0,R))$, $f(q)=\sum_{m=0}^{+\infty}q^{m}a_{m}$.
  1. a)
    $f$ is $\mathbb{C}_{I}$-preserving if and only if $a_{m}\in\mathbb{C}_{I}$ for all $m$.
  2. b)
    $f$ is slice preserving if and only if $a_{m}\in\mathbb{R}$ for all $m$.
Proof.
  1. a)
    If $a_{m}$ are all in $\mathbb{C}_{I}$, then clearly $f$ restricted to that slice has values in $\mathbb{C}_{I}$; on the other hand, for a slice preserving and regular function it is $f^{(m)}(q)\in\mathbb{C}_{I}$ for all $q$ in the $I$-slice, thus $a_{m}=(m!)^{-1}f^{(m)}(0)\in\mathbb{C}_{I}$.
  2. b)
    follows from a) since $\bigcap_{I\in\mathbb{S}}\mathbb{C}_{I}=\mathbb{R}$.∎
If $f\in\mathcal{SR}(B(0,R))$ is a slice preserving regular function and $g$ is slice regular on the same ball, then $f\ast g=fg$ and also $f\ast g=g\ast f$ by simply using power series expansion and the fact that the coefficients for $f$ are real. Furthermore, if $f,g$ preserve the same slice $\mathbb{C}_{I}$, then it is $f\ast g=g\ast f$.
If $f(q)=\sum_{m=0}^{+\infty}q^{m}a_{m}\in\mathcal{SR}(B(0,R))$, we define its regular conjugate function as $f^{c}(q):=\sum_{m=0}^{+\infty}q^{m}\overline{a}_{m}$ (of course, $f^{c}$ is still slice regular on $B(0,R)$) and its symmetrization as $f^{s}:=f\ast f^{c}=f^{c}\ast f$. One checks the coefficients of $f^{s}$ as a power series are real: indeed
$\left|\operatorname{Im}\left(\sum_{\ell=0}^{m}a_{m-\ell}\overline{a}_{\ell}\right)\right|=\frac{1}{2}\left|\sum_{\ell=0}^{m}a_{m-\ell}\overline{a}_{\ell}-\sum_{\ell=0}^{m}a_{\ell}\overline{a}_{m-\ell}\right|=0.$
[…]
$f^{-\ast}:=(f^{s})^{-1}\ast f^{c}=f^{c}\ast(f^{s})^{-1}=(f^{s})^{-1}f^{c}$
as a slice regular function on $B(0,R)\setminus\mathcal{Z}(f^{s})$: this has the property that $f^{-\ast}\ast f=f\ast f^{-\ast}=\operatorname{id}$. We call $f^{-\ast}$ the $\ast$-inverse of $f$. For any $q\in B(0,R)\setminus\mathcal{Z}(f^{s})$ it is $f^{-\ast}(q)=f^{s}(q)^{-1}f^{c}(q)=f(f^{c}(q)^{-1}qf^{c}(q))^{-1}$ by applying Proposition 1.10 to $f^{s}=f^{c}\ast f$. Analogously, if $g$ is another slice regular function in $B(0,R)$, then the (left) regular quotient $f^{-\ast}\ast g$ is defined and slice regular in $B(0,R)\setminus\mathcal{Z}(f^{s})$ and here it evaluates as
(3) $(f^{-\ast}\ast g)(q)=f^{-\ast}(q)\cdot g(f^{c}(q)^{-1}f^{s}(q)qf^{s}(q)^{-1}f^{c}(q))=(f^{-1}g)(f^{c}(q)^{-1}qf^{c}(q)),$
since by slice preservation of $f^{s}$ it is $f^{s}(q)qf^{s}(q)^{-1}=q$.
One may also define a right regular quotient $g\ast f^{-1}$, which is still slice regular in $B(0,R)\setminus\mathcal{Z}(f^{s})$; in general it is different from $f^{-\ast}\ast g$, but they coincide when $f$ and $g$ast-commute.
Remark 1.13.
The map $\mathcal{T}_{f}:q\mapsto f(q)^{-1}qf(q)$, which appears in the evaluation formula for the $\ast$-product, is not a slice regular map. However, if we restrict it to $B(0,R)\setminus\mathcal{Z}(f^{s})$, it is a real differentiable function which preserves the module and the real part: $|\mathcal{T}_{f}(q)|=|q|$ and $\operatorname{Re}(\mathcal{T}_{f}(q))=\operatorname{Re}q$. Moreover, the map $\mathcal{T}_{f^{c}}$ defined on $B(0,R)\setminus\mathcal{Z}(f^{s})$ by $q\mapsto f^{c}(q)^{-1}qf^{c}(q)$ is its inverse: $\mathcal{T}_{f}\circ\mathcal{T}_{f^{c}}=\mathcal{T}_{f^{c}}\circ\mathcal{T}_{f}=\operatorname{id}$ (see [16, Prop. 5.32]). So $\mathcal{T}_{f}$ is a diffeomorphism of $B(0,R)\setminus\mathcal{Z}(f^{s})$ onto itself.
1.2. Regular linear fractional and Möbius transformations
In the context of slice regularity theory, linear fractional transformations and Möbius transformation have been studied in [11, 22], starting from considerations on their ‘classical’ counterparts – without the $\ast$-product – already studied in [8]. We recall here some fact we need.
Definition 1.14.
A regular linear fractional transformation is a map of the form
$\mathcal{F}_{A}:q\mapsto(qc+d)^{-\ast}\ast(qa+b)$
associated to an invertible quaternionic matrix $A=\begin{pmatrix}a&c\\ b&d\end{pmatrix}\in\operatorname{GL}(2,\mathbb{H})$. A quaternionic $2\times 2$ matrix $A=\begin{pmatrix}a&c\\ b&d\end{pmatrix}$ is invertible if and only if its Dieudonné determinant (see [8]) is not zero:
$\operatorname{det}_{\mathbb{H}}A:=\sqrt{|a|^{2}|d|^{2}+|b|^{2}|c|^{2}-2\operatorname{Re}(b\overline{a}c\overline{d})}\neq 0.$
The map $\mathcal{F}_{A}$ belongs to the ring of quotients of $\mathcal{SR}(\mathbb{H})$ with respect to the $\ast$-product; it is slice regular on $\mathbb{H}$ if $c=0$ and on $\mathbb{H}\setminus\mathbb{S}_{dc^{-1}}$ otherwise, where $\mathbb{S}_{dc^{-1}}:=\{q:\operatorname{Re}q=\operatorname{Re}(dc^{-1}),|\operatorname{Im}q|=|\operatorname{Im}(dc^{-1})|\}=\mathcal{Z}((qc+d)^{s})$. By (3), one has $\mathcal{F}_{A}=F_{A}\circ\mathcal{T}$, where
$F_{A}(q)=(qc+d)^{-1}(qa+b),\qquad\mathcal{T}=\mathcal{T}_{q\overline{c}+\overline{d}}:q\mapsto(q\overline{c}+\overline{d})^{-1}\cdot q\cdot(q\overline{c}+\overline{d}).$
Remark 1.15.
It is clear that for all $\lambda\in\mathbb{R}\setminus\{0\}$, it is $\mathcal{F}_{\lambda A}=\mathcal{F}_{A}$, hence we can consider $A$ belonging to a class in the quotient group $\operatorname{PSL}(2,\mathbb{H})=\operatorname{GL}(2,\mathbb{H})/\{\lambda\mathbf{I}:\lambda\in\mathbb{R}\setminus\{0\}\}$. In particular, we may choose $\lambda=(\det_{\mathbb{H}}A)^{-1/2}$ and so we can assume $A\in\operatorname{SL}(2,\mathbb{H})=\{A\in\operatorname{GL}(2,\mathbb{H}):\det_{\mathbb{H}}A=1\}$.
Definition 1.16.
A regular linear fractional transformation $\mathcal{F}_{A}$ such that it is slice regular on the quaternionic unit ball $\mathbb{B}:=B(0,1)$ and such that $\mathcal{F}_{A}(\mathbb{B})\subseteq\mathbb{B}$ is said a (regular) Möbius transformation. They are characterized by the following result.
Proposition 1.17 ([22, Cor. 7.2]).
Let $\mathcal{F}_{A}$ be a regular linear fractional transformation for some $A\in\operatorname{GL}(2,\mathbb{H})$. The following are equivalent:
  1. a)
    $\mathcal{F}_{A}$ is a regular Möbius transformation.
  2. b)
[…]
  3. c)
    There exists (uniquely) $u\in\partial\mathbb{B}$ and $p\in\mathbb{B}$ such that
    $\mathcal{F}_{A}(q)=(1-q\overline{p})^{-\ast}\ast(q-p)u=(\overline{u}\ast q-p\overline{u})\ast(1-\overline{p}\ast q)^{-\ast}$
In particular, considering $u=1$ we get that for any $p\in\mathbb{B}$ we can associate a Möbius transformation
(4) $\mathcal{M}_{p}:=(1-q\overline{p})^{-\ast}\ast(q-p)=(q-p)\ast(1-q\overline{p})^{-\ast}$
which is associated to an Hermitian matrix
$\frac{1}{\sqrt{1-|p|^{2}}}\begin{pmatrix}1&-\overline{p}\\ -p&1\end{pmatrix}\in\operatorname{Sp}(1,1).$
Evaluating $\mathcal{M}_{p}$ at any point $q$ in the unit ball gives $\mathcal{M}_{p}(q)=(M_{p}\circ\mathcal{T}_{p})(q)$ where $M_{p}(q)$ is the ‘classical’ Möbius map $(1-q\overline{p})^{-1}(q-p)$ and $\mathcal{T}_{p}(q)=(1-qp)^{-1}q(1-qp)$. Both these maps $M_{p}$ and $\mathcal{T}_{p}$ are bijective from $\mathbb{B}$ to itself, with inverse maps
$M_{p}^{-1}=M_{-p}$ $:s\mapsto(1+s\overline{p})^{-1}(s+p),$
$\mathcal{T}_{p}^{-1}=\mathcal{T}_{\overline{p}}$ $:s\mapsto(1-s\overline{p})^{-1}s(1-s\overline{p}).$
Since right multiplication by a unimodular constant is an invertible operation too, we have proved that:
Proposition 1.18.
All regular Möbius transformations are bijective from $\mathbb{B}$ onto itself.
Moreover, we remark that any point outside the unit ball in which a given regular Möbius transformation is slice regular is mapped outside the unit ball.
A relevant fact of these maps is that – if they are identified with the multiplicative group $\operatorname{Sp}(1,1)$ – they act on the space slice regular functions mapping the unit ball to itself $\mathcal{SR(\mathbb{B},\mathbb{B})}=\{f\in\mathcal{SR}(\mathbb{B}):f(\mathbb{B})\subseteq\mathbb{B}\}$.
Proposition 1.19 ([11, 22]).
Let $\mathcal{F}_{A}$ be a regular Möbius transformation and let us assume that the associated matrix is $A=(1-|p|^{2})^{-1/2}\begin{pmatrix}u&-\overline{p}\\ -pu&1\end{pmatrix}\in\operatorname{Sp}(1,1)$ for some $u\in\partial\mathbb{B}$ and $p\in\mathbb{B}$, i.e., $\mathcal{F}_{A}=\mathcal{M}_{p}u$. If $f\in\mathcal{SR(\mathbb{B},\mathbb{B})}$, then the function
$f.A:=(1-f\overline{p})^{-\ast}\ast(f-p)u$
[…]
Alternatively, we can define a left action of $\operatorname{Sp}(1,1)$ over $\mathcal{SR(\mathbb{B},\mathbb{B})}$ by
[…]
An easy consequence of what stated above is that any regular Möbius transformation is bijective.
[…]
$(\mathcal{M}_{p}\bullet f)(q)=(1-f\overline{p})^{-\ast}\ast(f-p)\mid_{q}=(1-f(\widetilde{\mathcal{T}}(q))\overline{p})^{-1}(f(\widetilde{\mathcal{T}}(q))-p)=(M_{p}\circ f\circ\widetilde{\mathcal{T}})(q)$
where $\widetilde{\mathcal{T}}(q)=\varphi_{f}(q)^{-1}\cdot q\cdot\varphi_{f}(q)$, $\varphi_{f}(q)=(1-f\ast\overline{p})^{c}(q)=1-(p\ast f^{c})(q)$.
1.3. Regular Blaschke products
We now introduce the analogous notion of a complex Blaschke product in the framework of quaternionic slice regularity, which already appeared in [3, 4, 18].
Definition 1.20.
A (slice) regular (finite) Blaschke product is a function $f$ over $\mathbb{B}$ which can be written in the form
(5) $f=\mathcal{M}_{s_{1}}\ast\mathcal{M}_{s_{2}}\ast\ldots\ast\mathcal{M}_{s_{\kappa}}u,\qquad(\kappa\geq 1)$
where $u\in\partial\mathbb{B}$ is a unimodular constant and $s_{1},\ldots,s_{\kappa}\in\mathbb{B}$.
If $f$ is of this form, it clearly maps the unit ball to itself, it is slice regular, it extends continuously to the unit sphere $\partial\mathbb{B}$ which is $f$-invariant.
The zero set of this function, and in general of any slice regular function, can be described as a combination of (eventually repeated) $2$-spheres $\mathbb{S}_{p_{1}},\ldots,\mathbb{S}_{p_{m}}$ where $\mathbb{S}_{p_{\ell}}=\{q:\operatorname{Re}q=\operatorname{Re}(p_{\ell}),|q|=|p_{\ell}|\}$ and points $q_{1},\ldots,q_{n}$, which could be not distinct or not disjoint from the spheres, such that the total multiplicity is $n+2m=\kappa$. For these definitions and the general description of the zero sets of slice regular functions we refer to Chapter 3 of [16].
This last remark says that, although the expression (5) is not unique, the number $\kappa$ is uniquely determined and it is called the degree of the Blaschke product $f$. Blaschke products of degree $1$, $f=\mathcal{M}_{q_{1}}u$, are exactly the regular Möbius transformations.
Now, the properties listed above give a complete characterization of slice regular Blaschke products.
Lemma 1.21.
Let $f:\mathbb{B}\longrightarrow\mathbb{B}$ be a slice regular function such that
  1. a)
    it extends continuously up to $\partial\mathbb{B}$ and $|f|\equiv 1$ on $\partial\mathbb{B}$;
  2. b)
    its set of zeros is given by a finite number of points $q_{1},\ldots,q_{n}$ and spheres $\mathbb{S}_{p_{1}},\ldots,\mathbb{S}_{p_{m}}$ (counted with their respective multiplicity), with $\kappa=n+2m$.
Then $f=Bu$, where $B$ is a regular Blaschke product of degree $\kappa$ with the same zero set and $u$ is a unimodular quaternionic constant.
Proof.
Let $B$ be a Blaschke product with zero set given by the points $q_{1},\ldots,q_{n}$ and the spheres $\mathbb{S}_{p_{1}},\ldots,\mathbb{S}_{p_{m}}$ with correspondent multiplicity. Let $f$ be as in the assumption. Then by a factorization argument one shows that both $f^{-\ast}\ast B$ and $B^{-\ast}\ast f$ are regular functions over $\mathbb{B}$, with continuous extension to the boundary. Moreover $|f^{-\ast}\ast B|=|B^{-\ast}\ast f|=1$ on the whole $\partial\mathbb{B}$. Hence by the Maximum Modulus Principle [16, Thm 7.1] it is
$|B^{-\ast}\ast f|\leq 1,\qquad|f^{-\ast}\ast B|\leq 1\qquad\text{on }\mathbb{B}.$
Let $T_{f}(q)=f^{c}(q)^{-1}qf^{c}(q)$ and $T_{B}(q)$ the same with $f$ replaced by $B$. Let $q_{0}$ be a point in the unit ball which is not a zero of $f^{s}$ (and thus of $B^{s}$), i.e., $q_{0}\in\mathbb{B}\setminus\left[\left(\bigcup_{h=1}^{n}\mathbb{S}_{q_{h}}\right)\cup\left(\bigcup_{\ell=1}^{m}\mathbb{S}_{p_{\ell}}\right)\right]$.  [^2] If $r\in\mathbb{R}$, then the 2-sphere $\mathbb{S}_{r}=\{r\}$ collapses to a single real point. In this set the maps above are real diffeomorphisms (hence invertible). So we can define $\widetilde{q_{0}}=T_{B}^{-1}(q_{0})$, $\widehat{q_{0}}=T_{f}^{-1}(q_{0})$ and we have:
$(B^{-\ast}\ast f)(\widetilde{q_{0}})=B(q_{0})^{-1}f(q_{0})=:u\qquad\Longrightarrow\qquad|u|\leq 1;$
$(f^{-\ast}\ast B)(\widehat{q_{0}})=f(q_{0})^{-1}B(q_{0})=u^{-1}\qquad\Longrightarrow\qquad|u|\geq 1.$
Therefore there exists a point $q_{0}\in\mathbb{B}$ such that $B^{-\ast}\ast f$ evaluated at $q_{0}$ gives $u$, which has modulus one, hence by the Maximum Modulus Principle it is $B^{-\ast}\ast f\equiv u$, that is to say $f=Bu$. ∎
Proposition 1.22.
Let $B$ be a regular Blaschke product of degree $\kappa$. If $\mathcal{M}_{q_{0}}$ is the regular Möbius transformation associated to a point $q_{0}\in\mathbb{B}$, then $\mathcal{M}_{q_{0}}\bullet B$ is again a regular Blaschke product with the same degree.
Conversely, if $f\in\mathcal{SR}(\mathbb{B})$ is such that $\mathcal{M}_{q_{0}}\bullet f$ is a Blaschke product of degree $\kappa$, then $f$ is a Blaschke product of degree $\kappa$.
Proof.
We have that $\mathcal{M}_{q_{0}}\bullet B(q)=(B(q)-q_{0})\ast(1-\overline{q_{0}}\ast B(q))^{-\ast}$ is a slice regular function which clearly extends continuously to the boundary. Moreover it is not difficult to check that $|(\mathcal{M}_{q_{0}}\bullet B)(q)|\rightarrow 1$ as $q$ goes to any point of $\partial\mathbb{B}$. In order to conclude it remains to consider its zero set. So we need to consider the fiber $B^{-1}(q_{0})$.
Let us assume $B=\mathcal{M}_{s_{1}}\ast\cdots\ast\mathcal{M}_{s_{\kappa}}$ (multiplying for a unimodular constant will not change the argument). The equation $B(q)=q_{0}$, where $q_{0}\in\mathbb{B}$, will admit only solutions of modulus less than one, as otherwise there would be a Möbius factor whose value in a point outside $\mathbb{B}$ is of norm less than one, and that is absurd.
If $\kappa=1$, it is clear that there is a unique solution of $B(q)=q_{0}$:
$B(q)=q_{0}\iff q-s_{1}=(1-q\overline{s_{1}})q_{0}\iff q=(s_{1}+q_{0})(1+\overline{s_{1}}q_{0})^{-1}.$
If $\kappa=2$, $B(q)=q_{0}$ can be rewritten as
$(1-q\overline{s_{1}})^{-\ast}\ast(q-s_{1})\ast(q-s_{2})\ast(1-q\overline{s_{2}})^{-\ast}=q_{0},$
hence by left $\ast$-multiplying by $(1-q\overline{s_{1}})$ and right $\ast$-multiplying by $(1-q\overline{sp_{2}})$, then
$(q-s_{1})\ast(q-s_{2})=(1-q\overline{s_{1}})q_{0}\ast(1-q\overline{s_{2}}),$
thus we get the regular polynomial equation
$q^{2}(1-\overline{s_{1}}q_{0}\overline{s_{2}})+q(\overline{s_{1}}q_{0}+q_{0}\overline{s_{2}}-s_{1}-s_{2})+s_{1}s_{2}-q_{0}=0.$
The leading coefficient is not zero, so we reduce ourselves to consider the zero set of a slice regular polynomial of degree $2$, which has total multiplicity $2$.
Now, when $\kappa\geq 3$, we need to be more careful. Indeed now some of the ‘star-denominators’ are not the left or the right $\ast$-factor in the expression of $B(q)$, so we need to find a way to shift them. With this purpose, we observe that
$(1-q\alpha)^{-\ast}\ast(q-\beta)=(q-\beta\gamma)\ast(\gamma-q\alpha)^{-\ast}$
where $\gamma=(1-\alpha\beta)^{-1}(1-\beta\alpha)\in\partial\mathbb{B}$. Indeed:
$(1-q\alpha)\ast(q-\beta\gamma)=q^{2}(-\alpha)+q(1+\alpha\beta\gamma)-\beta\gamma,$
$(q-\beta)\ast(\gamma-q\alpha)=q^{2}(-\alpha)+q(\gamma+\beta\alpha)-\beta\gamma,$
$1+\alpha\beta\gamma=\gamma+\beta\alpha\Leftrightarrow(1-\alpha\beta)\gamma=1-\beta\alpha.$
Let us remark that $|\gamma|=1$. Using this trick and then proceeding as in the $\kappa=2$ case, one gets that the equation $B(q)=q_{0}$ is equivalent to $P_{\kappa}(q)=0$, where $P_{\kappa}$ is a slice regular polynomial of degree $\kappa$, hence the set of its roots has total multiplicity $\kappa=n+2m$, where $m$ is number of spherical zeros and $n$ number of non-spherical zeros (counted with their respective multiplicity).
For the converse part, it is enough to observe that if $\mathcal{M}_{q_{0}}\bullet f$ is a Blaschke product, then we can take $\mathcal{M}_{-q_{0}}$: as the associated matrices in $\operatorname{Sp}(1,1)$ are such that their product is the identity matrix $\mathbb{I}$, then $\mathcal{M}_{-q_{0}}\bullet(\mathcal{M}_{q_{0}}\bullet f)=f$. Thus $f$ is a Blaschke product too. ∎
In the particular case $\kappa=2$, we have also:
Proposition 1.23.
A finite Blaschke product $B$ is of degree two if and only if it exists a Möbius transformation $\mathcal{M}_{p}$ and a point $q_{c}\in\mathbb{B}$ such that $(\mathcal{M}_{p}\bullet B)u=\mathcal{M}_{q_{c}}^{\ast 2}=\mathcal{M}_{q_{c}}\ast\mathcal{M}_{q_{c}}$ for some $u\in\partial\mathbb{B}$.
Proof.
The if part easily follows by the previous result.
If we assume that $B=\mathcal{M}_{q_{1}}\ast\mathcal{M}_{q_{2}}$, without losing generality, then
(6) $\partial_{C}B(q)=(1-q\overline{q_{1}})^{-\ast 2}\ast[(1-|q_{1}|^{2})((q-q_{2})\ast(1-q\overline{q_{2}}))+(1-|q_{2}|^{2})((1-q\overline{q_{1}})\ast(q-q_{1}))]\ast(1-q\overline{q_{2}})^{-\ast 2}$
The term within the square brackets in (6) writes as a regular polynomial $q^{2}a_{2}+2qa_{1}+a_{0}$, where
$a_{0}$ $=$ $-(1-|q_{1}|^{2})q_{2}-(1-|q_{2}|^{2})q_{1}$
$a_{1}$ $=$ $1-|q_{1}|^{2}|q_{2}|^{2}\in\mathbb{R},$
$a_{2}$ $=$ $\overline{a_{0}}=-(1-|q_{1}|^{2})\overline{q_{2}}-(1-|q_{2}|^{2})\overline{q_{1}}.$
When $q_{1}=-q_{2}$, this polynomial becomes $2q(1-|q_{1}|^{4})$ which has a unique root $q=0$. Otherwise, we have a polynomial of degree two for which $q=0$ is not a root. A priori, we get that the set of zeros is non-empty in $\mathbb{H}$ (in particular, of total multiplicity two). Now it is clear that the coefficients above lie all in the same slice $\mathbb{C}_{I}$ for some $I\in\mathbb{S}$, so we may solve the equation as in the complex case and the roots will lie on the same slice $\mathbb{C}_{I}$ as the coefficients. These roots are also critical points of the restriction $B\mid_{\mathbb{C}_{I}}$ which is a complex Blaschke product by identifying $\mathbb{C}$ with $\mathbb{C}_{I}$. We know that in the complex case a Blaschke product of degree two admits a unique critical point inside the unit disk $\mathbb{D}\cong\mathbb{B}\cap\mathbb{C}_{I}$. We conclude that there exists $q_{c}\in\mathbb{B}$ such that $(\partial_{C}B)(q_{c})=0$. Let $p_{c}=B(q_{c})$. Now by choosing $\mathcal{M}=\mathcal{M}_{p_{c}}$, then $\mathcal{M}_{p_{c}}\bullet B$ is a Blaschke product of degree two which vanishes at $q_{c}$. Now
$\mathcal{M}_{p_{c}}\bullet B$ $=$ $(B-p_{c})\ast(1-\overline{p_{c}}\ast B)^{-\ast},$
$\partial_{C}(\mathcal{M}_{p_{c}}\bullet B)$ $=$ $\partial_{C}B\ast(1-\overline{p_{c}}\ast B)^{-\ast}+(B-p_{c})\ast\partial_{C}((1-\overline{p_{c}}\ast B)^{-\ast}),$
hence $(\partial_{C}(\mathcal{M}_{p_{c}}\bullet B))(q_{c})=0$ as $B(q_{c})-p_{c}=0$ and $(\partial_{C}B)(q_{c})=0$. We get that the zero set of $\mathcal{M}_{p_{c}}\bullet B$ is given by $q_{c}$ counted twice, so it coincides with $\mathcal{M}_{q_{c}}^{\ast 2}$ up to right multiplication by a unimodular constant. ∎
We conclude this section by recalling that – although we will not consider them in the following – it is also possible to define regular infinite Blaschke products. We refer to [3, 18] for their definition.
2. The hyperbolic difference quotient and the three-points Schwarz–Pick Lemma
We start by recalling how to define a difference quotient in the context of the theory of slice regularity.
Definition 2.1 ([10, Rmk. 1.6]).
Let $f:\mathbb{B}\longrightarrow\mathbb{B}$ be a slice regular function. Fix a point $p\in\mathbb{B}$. Then the regular difference quotient of $f$ at $p$ is the function
$\mathcal{R}_{f,p}(q):=(q-p)^{-\ast}\ast(f(q)-f(p)).$
At a first look, since the term $(q-p)^{-\ast}$ is well defined only for $q\in\mathbb{B}$ not in the sphere $\mathbb{S}_{p}=\{q:q^{2}-q\cdot 2\operatorname{Re}p+|p|^{2}=0\}$, the domain of $\mathcal{R}_{f,p}$ is only $\mathbb{B}\setminus\mathbb{S}_{p}$. But the singularity at $\mathbb{S}_{p}$ is removable: indeed, by applying [16, Prop. 3.18] to $f-f(p)$, there exists a unique slice regular function $g:\mathbb{B}\longrightarrow\mathbb{B}$ for which $f(q)=f(p)+(q-p)\ast g(q)$, and it is clear that $g$ coincides with $\mathcal{R}_{f,p}$ in $\mathbb{H}\setminus\mathbb{S}_{p}$. The value in $p$ of this function coincides with the regular derivative $\partial_{C}f(p)$, whereas its value in $\overline{p}$ coincides by the very definition with the spherical derivative in $p$: $\partial_{S}f(p):=(2\operatorname{Im}p)^{-1}(f(p)-f(\overline{p}))$.
Definition 2.2.
Let $f:\mathbb{B}\longrightarrow\mathbb{B}$ be a slice regular function. Fix a point $p\in\mathbb{B}$. We define the hyperbolic difference quotient at $p$
$f^{\star}_{p}(q):=(1-q\overline{p})\ast\mathcal{R}_{f,p}\ast(1-\overline{f(p)}\ast f(q))^{-\ast}.$
Then $f_{p}^{\star}:\mathbb{B}\longrightarrow\mathbb{H}$ is a slice regular function as a $\ast$-product of regular functions: the last term is not singular since $f$ has values in the unit ball. In particular, its value in $q=p$ is called the hyperbolic derivative of $f$ in $p$ and it is denoted by $f^{h}(p)$. By evaluating the $\ast$-products, it is related to the regular derivative by the formula
(7) $f^{h}(p)=f_{p}^{\star}(p)=(1-|p|^{2})\partial_{C}f(p)(1-\overline{f(p)}f(p_{0}))^{-1},$
for some $p_{0}\in\mathbb{S}_{p}$ depending on $p$ and on the function $f$ itself. Note that $f_{p}^{\star}$ can also be written as
$\boxed{f_{p}^{\star}=\mathcal{M}_{p}^{-\ast}\ast(\mathcal{M}_{f(p)}\bullet f)}$
where $\mathcal{M}_{p}$ is the regular Möbius transformation (4) and by $\mathcal{M}_{f(p)}\bullet f$ we mean the (right) action on $f$ of the matrix $\frac{1}{\sqrt{1-|f(p)|^{2}}}\begin{pmatrix}1&-\overline{f(p)}\\ -f(p)&1\end{pmatrix}$ given by Proposition 1.19, namely
(8) $\mathcal{M}_{f(p)}\bullet f=(f-f(p))\ast(1-\overline{f(p)}\ast f)^{-\ast}=(1-f\cdot\overline{f(p)})^{-\ast}\ast(f-f(p)).$
2.1. The quaternionic Schwarz–Pick Lemma
The following results have been obtained by Bisi and Stoppato in [10].
Proposition 2.3 ([10, Thm. 3.2]).
$f:\mathbb{B}\longrightarrow\mathbb{B}$ be a slice regular function. Let $p\in\mathbb{B}$ such that $f(p)=0$. Then for all $q\in\mathbb{B}$ it is
(9) $|(\mathcal{M}_{p}^{-\ast}\ast f)(q)|\leq 1.$
The inequality is strict unless $f=\mathcal{M}_{p}u$ for some $u\in\partial\mathbb{B}$, i.e., $f$ is a Möbius transformation vanishing at $p$.
Theorem 2.4 (Quaternionic Schwarz–Pick Lemma [10, Thm. 3.7]).
Let $f:\mathbb{B}\longrightarrow\mathbb{B}$ be a slice regular function. Let $p\in\mathbb{B}$. Then for all $q\in\mathbb{B}$
(10) $|(\mathcal{M}_{f(p)}\bullet f)(q)|\leq|\mathcal{M}_{p}(q)|,$
$|\mathcal{R}_{f,p}\ast(1-\overline{f(p)}\ast f)^{-\ast}|_{\mid_{q}}\leq|(1-\overline{p}\ast q)^{-\ast}|.$
Moreover, for any $p\in\mathbb{B}$
$|\partial_{C}f\ast(1-\overline{f(p)}\ast f)^{-\ast}|_{\mid_{p}}\leq\dfrac{1}{1-|p|^{2}},$
$\dfrac{|\partial_{S}f(p)|}{|1-f^{s}(p)|}\leq\dfrac{1}{|1-\overline{p}^{2}|}.$
If $f$ is a regular Möbius transformation, i.e. of the form $\mathcal{M}_{p}u$ where $|u|=1$, then all formulas above are equalities. Otherwise, the inequality (10) is strict for any $q\in\mathbb{B}\setminus\{p\}$ while the others are strict on the whole unit ball.
Remark 2.5.
Recall that, if we denote $M_{p}(q)=(1-q\overline{p})^{-1}(q-p)$, then $\mathcal{M}_{p}(q)$ evaluates as $M_{p}((1-qp)^{-1}q(1-qp))$. So the right hand side of (10) can be interpreted, once $q\in\mathbb{B}$ is fixed, as the (classical) pseudo-hyperbolic distance between $p$ and $\mathcal{T}_{p}(q):=(1-qp)^{-1}q(1-qp)$. On the other hand,
$(\mathcal{M}_{f(p)}\bullet f)(q)=(M_{f(p)}\circ f)(\widetilde{\mathcal{T}}_{f,p}(q)),$
where
$\widetilde{\mathcal{T}}_{f,p}(q):=\phi(q)^{-1}\cdot q\cdot\phi(q),\quad\phi(q)=(1-f\ast\overline{f(p)})^{c}(q).$
The maps $\mathcal{T}_{p}$ and $\widetilde{\mathcal{T}}_{f,p}$ are not regular maps, however they are real diffeomorphisms leaving the spheres of type $\mathbb{S}_{p}$ invariant, where defined: indeed it is straightforward to check that they preserve the real part and the modulus of any quaternion in the unit ball.
If $\rho_{\mathbb{B}}$ denotes the classical pseudo-hyperbolic distance $\rho_{\mathbb{B}}(p,q):=|M_{p}(q)|$, then we have
(11) $\rho_{\mathbb{B}}(f(p),f(\widetilde{\mathcal{T}}_{f,p}(q)))\leq\rho_{\mathbb{B}}(p,\mathcal{T}_{p}(q)).$
The result is therefore not in general as in the complex case, where in the left hand side we just have that the pullback of the pseudo-hyperbolic distance with respect to an holomorphic function is less or equal than the pseudo-hyperbolic distance itself. Here the two maps $\mathcal{T}_{p}$ and $\widetilde{\mathcal{T}}_{f,p}$ appearing in the two sides are different.
We could also consider the regular pseudo-hyperbolic distance $\rho^{\ast}_{\mathbb{B}}(p,q):=|\mathcal{M}_{p}(q)|$. As stated in [5], it is $\rho^{\ast}_{\mathbb{B}}(p,q)=\rho^{\ast}_{\mathbb{B}}(q,p)$ and it is the distance associated to the reproducing kernel Hilbert space $H^{2}(\mathbb{B})$ with kernel $k_{p}(q)=(1-q\overline{p})^{-\ast}$. Then $\rho^{\ast}_{\mathbb{B}}(p,q)$ is exactly the right hand side of equation (2.7) and we have the relation
$\rho^{\ast}_{\mathbb{B}}(p,q)=\rho_{\mathbb{B}}(p,\mathcal{T}_{p}(q)).$
Now $\mathcal{T}_{p}$ is a real diffeomorphism on $\mathbb{B}$ with inverse $\mathcal{T}_{p}^{-1}=\mathcal{T}_{\overline{p}}:q\mapsto(1-q\overline{p})^{-1}q(1-q\overline{p})$, so $\rho_{\mathbb{B}}(p,w)=\rho^{\ast}_{\mathbb{B}}(p,\mathcal{T}_{p}^{-1}(w))=\rho^{\ast}_{\mathbb{B}}(p,\mathcal{T}_{\overline{p}}(w))$. Hence (10) in terms of $\rho^{\ast}_{\mathbb{B}}$ becomes
$\rho^{\ast}_{\mathbb{B}}\left(f(p),\mathcal{T}_{\overline{f(p)}}(f(\widetilde{\mathcal{T}}_{f,p}(q)))\right)\leq\rho^{\ast}_{\mathbb{B}}(p,q).$
In this way, it is clear that the use of the distance $\rho_{\mathbb{B}}^{\ast}$ does not give a simpler interpretation of Theorem 2.7. Analogous considerations can be made for the inequality (10) of Theorem 2.4.
In the following we will also use the Poincaré distance:
$\delta_{\mathbb{B}}(p,q):=\tanh^{-1}\rho_{\mathbb{B}}(p,q)=\frac{1}{2}\log\frac{1+\rho_{\mathbb{B}}(p,q)}{1-\rho_{\mathbb{B}}(p,q)}.$
By monotonicity of the function $\tanh$, (11) still holds if we replace the pseudo-hyperbolic distance with the Poincaré distance:
(12) $\delta_{\mathbb{B}}(f(p),f(\widetilde{\mathcal{T}}_{f,p}(q)))\leq\delta_{\mathbb{B}}(p,\mathcal{T}_{p}(q)).$
This Poincaré metric has been first introduced in [8] for the quaternionic setting. See also [9] for a comparison with the Kobayashi distance of $\mathbb{B}$ as a subset of $\mathbb{C}^{2}$.
2.2. Three-points quaternionic Schwarz–Pick Lemma
Now, let us return to the hyperbolic difference quotient $f^{\star}_{p}$: we can apply the previous results obtaining a Schwarz–Pick type statement which now involves three points of the unit ball.
Theorem 2.6.
Let $f:\mathbb{B}\longrightarrow\mathbb{B}$ be slice regular and $p\in\mathbb{B}$. For any $q\in\mathbb{B}$ it is $|f^{\star}_{p}(q)|\leq 1$. The strict inequality holds, except for the case when $f$ is a regular Möbius transformation (times a unimodular constant). In the latter situation, it is $f^{\star}_{p}\equiv u\in\partial\mathbb{B}$.
Proof.
The result follows by (9) where $f$ is replaced by the function $g:=\mathcal{M}_{f(p)}\bullet f$, which has a zero at the point $p$, so $\mathcal{M}_{g(p)}\bullet g=\mathcal{M}_{0}\bullet g=g$. ∎
Theorem 2.7 (Three-points Schwarz–Pick Lemma).
Suppose $f:\mathbb{B}\longrightarrow\mathbb{B}$ is slice regular but not a regular Möbius transformation. Let $p,s\in\mathbb{B}$. Then for any $q\in\mathbb{B}$
$|(\mathcal{M}_{f_{p}^{\star}(s)}\bullet f^{\star}_{p})(q)|\leq|\mathcal{M}_{s}(q)|.$
Proof.
Just apply (10) to the function $f_{p}^{\star}$. ∎
To consider when equality holds in the theorem above, we observe the following:
Proposition 2.8.
Suppose $f:\mathbb{B}\longrightarrow\mathbb{B}$ is slice regular but not a regular Möbius transformation and fix $p\in\mathbb{B}$. Then $f^{\star}_{p}$ is a regular Möbius transformation (up to right multiplication by unimodular constants) if and only if $f$ is a regular Blaschke product of degree $2$.
Proof.
By the very definition of $f_{p}^{\star}$, it is
$\mathcal{M}_{p}\ast f_{p}^{\star}=\mathcal{M}_{f(p)}\bullet f.$
Then it is clear that if $f_{p}^{\star}$ is a regular Möbius transformation, then $\mathcal{M}_{f(p)}\bullet f$ is a regular Blaschke product of degree two. Hence $f$ is also a regular Blaschke product of degree $2$. On the other side, if $f$ is a regular Blaschke product of degree $2$, then the right hand side is a regular Blaschke product vanishing at $p$, hence it writes down as $\mathcal{M}_{p}\ast\mathcal{M}_{p^{\prime}}u$ for some $p^{\prime}\in\mathbb{B}$, implying that $f_{p}^{\star}=\mathcal{M}_{p^{\prime}}u$. ∎
As a direct consequence one gets that:
Theorem 2.9.
Equality holds in (2.7) for some $q\in\mathbb{B},q\neq s$ if and only if the function $f$ is a regular Blaschke product of degree $2$. In this case, equality holds for all $q\in\mathbb{B}$.
Remark 2.10.
In the proof of Theorem 2.4 given in [10], it is claimed that $\partial_{S}(\mathcal{M}_{f(p)}\bullet f)(p)=(1-\overline{f^{s}(p)})^{-1}\partial_{S}f(p)$. From this we can get a formula for $f^{\star}_{p}$ evaluated in $\overline{p}$:
(13) $f_{p}^{\star}(\overline{p})=(1-\overline{p}^{2})\mathcal{R}_{(\mathcal{M}_{f(p)}\bullet f),p}(\overline{p})=(1-\overline{p}^{2})(1-\overline{f^{s}(p)})^{-1}\partial_{S}f(p).$
3. Some consequences. Estimates in the case of $\boldsymbol{f(0)=0}$
Lemma 3.1.
Let $f:\mathbb{B}\longrightarrow\mathbb{B}$ be slice regular and such that $f(0)=0$. Then for any $q_{0}\in\mathbb{B}$ it is
$f^{\star}_{0}(q_{0})=f^{\star}_{q_{0}}(0)=q_{0}^{-1}f(q_{0}).$
Proof.
Since $\mathcal{M}_{0}$ is the identity function $q\mapsto q$ and $\mathcal{M}_{0}\bullet f=f$, it is clear that $f^{\star}_{0}(q_{0})=(\mathcal{M}_{0}^{-\ast}\ast(\mathcal{M}_{0}\bullet f))(q_{0})=q_{0}^{-1}f(q_{0})$. On the other side, since 00 is invariant by the rotations appearing by evaluating the $\ast$-products punctually, then
$f^{\star}_{q_{0}}(0)=(\mathcal{M}_{q_{0}}^{-\ast}\ast(\mathcal{M}_{f(q_{0})}\bullet f))(0)=M_{q_{0}}(0)^{-1}M_{f(q_{0})}(0)=(-q_{0})^{-1}(-f(q_{0}))=q_{0}^{-1}f(q_{0}).\qed$
More generally, we have the following
Lemma 3.2.
Let $f:\mathbb{B}\longrightarrow\mathbb{B}$ be slice regular. Whenever $r,s\in(-1,1)$ are real, it is $\left|f^{\star}_{r}(s)\right|=\left|f^{\star}_{s}(r)\right|$.
Proof.
Using the fact that $\mathcal{M}_{r}=M_{r}$ is a slice preserving function and that $\mathcal{M}_{f(r)}\bullet f$ coincides with $M_{f(r)}\circ f$ on the real interval $(-1,1)$, one gets
$f^{\star}_{r}(s)=(\mathcal{M}_{r}^{-\ast}\ast(\mathcal{M}_{f(r)}\bullet f))(s)=M_{r}(s)^{-1}M_{f(r)}(f(s)).$
Analogously,
$f^{\star}_{s}(r)=(\mathcal{M}_{s}^{-\ast}\ast(\mathcal{M}_{f(s)}\bullet f))(r)=M_{s}(r)^{-1}M_{f(s)}(f(r)).$
Now one concludes by recalling that $|M_{p}(q)|=|M_{q}(p)|$ for any $q,p\in\mathbb{B}$, as
$M_{p}(q)=-(1-q\overline{p})^{-1}(1-p\overline{q})M_{q}(p)$
and the factor $-(1-q\overline{p})^{-1}(1-p\overline{q})$ has clearly modulus one. ∎
Lemma 3.2 is important to get the following ‘four-point’ result (see [7, Cor 3.3]), where we have to assume that two of the points are real. Here we consider the Poincaré distance $\delta_{\mathbb{B}}$.
Corollary 3.3.
Let $r,s\in(-1,1)$ and $q,p\in\mathbb{B}$. For $f:\mathbb{B}\longrightarrow\mathbb{B}$ slice regular function which is not a regular Möbius transformation, we have
$\delta_{\mathbb{B}}(0,f^{\star}_{r}(q))\leq\delta_{\mathbb{B}}(0,f^{\star}_{s}(p))+\delta_{\mathbb{B}}(p,r)+\delta_{\mathbb{B}}(q,s).$
Proof.
By the triangle inequality and (12) with $p=0,q=s$ and $f$ replaced by $f^{\star}_{r}$, it is
$\delta_{\mathbb{B}}(0,f^{\star}_{r}(q))\leq\delta_{\mathbb{B}}(0,f^{\star}_{r}(s))+\delta_{\mathbb{B}}(f^{\star}_{r}(q),f^{\star}_{r}(s))\leq\delta_{\mathbb{B}}(0,f^{\star}_{r}(s))+\delta_{\mathbb{B}}(q,s).$
By Lemma 3.2 $|f^{\star}_{r}(s)|=|f^{\star}_{s}(r)|$, so their Poincaré distance from the origin is the same. Thus
$\delta_{\mathbb{B}}(0,f^{\star}_{r}(q))\leq\delta_{\mathbb{B}}(0,f^{\star}_{s}(r))+\delta_{\mathbb{B}}(q,s).$
Applying, as above, triangle inequality and (12) to the first term on the right hand side, one gets
$\delta_{\mathbb{B}}(0,f^{\star}_{r}(q))\leq\delta_{\mathbb{B}}(0,f^{\star}_{s}(p))+\delta_{\mathbb{B}}(p,r)+\delta_{\mathbb{B}}(q,s).\qed$
Lemma 3.4.
Let $c_{0}\in\mathbb{B}$ and $10<r_{0}<$ 1. Let us consider the pseudo-hyperbolic ball $B_{\rho}(c_{0},r_{0}):=\{q\in\mathbb{B}:\rho_{\mathbb{B}}(q,c_{0})<r_{0}\}$. This coincides with the Euclidean ball $B(c_{1},r_{1}):=\{q\in\mathbb{H}:|q-c_{1}|<r_{1}\}$ where
$\begin{cases}c_{1}=\dfrac{1-r_{0}^{2}}{1-|c_{0}|^{2}r_{0}^{2}}c_{0},\\ r_{1}=\dfrac{1-|c_{0}|^{2}}{1-|c_{0}|^{2}r_{0}^{2}}r_{0}.\end{cases}$
Proof.
By observing that the pseudo-hyperbolic distance is invariant by multiplication with a unimodular constant, we may assume that $c_{0}=|c_{0}|\in[0,1)$ and then deduce the general case by applying the inverse rotation. With this assumption, for any $I\in\mathbb{S}$ we observe that $B_{\rho}(c_{0},r_{0})\cap\mathbb{B}_{I}$ is an open disk which is symmetric with respect to the real line, described by the equation [^3] Symmetry w.r.t. the real line follows by invariance with respect to conjugation $q\mapsto\overline{q}$ $|q-c_{0}|<r_{0}|1-c_{0}q|$ Hence, to find the (Euclidean) center and radius of such disk, we compute the intersections of the boundary with the real line:
(14) $x-c_{0}=\pm r_{0}(1-c_{0}x)\implies x_{\pm}=\dfrac{c_{0}\pm r_{0}}{1\pm c_{0}r_{0}}.$
Hence $B_{\rho}(c_{0},r_{0})\cap\mathbb{B}_{I}$ corresponds to a disk along the slice of center
$c_{1}=\frac{1}{2}(x_{+}+x_{-})=\frac{1}{2}\dfrac{(c_{0}+r_{0})(1-c_{0}r_{0})+(c_{0}-r_{0})(1+c_{0}r_{0})}{1-c_{0}^{2}r_{0}^{2}}=\dfrac{c_{0}(1-r_{0}^{2})}{1-c_{0}^{2}r_{0}^{2}}$
and radius
$r_{1}=\frac{1}{2}(x_{+}-x_{-})=\frac{1}{2}\dfrac{(c_{0}+r_{0})(1-c_{0}r_{0})-(c_{0}-r_{0})(1+c_{0}r_{0})}{1-c_{0}^{2}r_{0}^{2}}=\dfrac{r_{0}(1-c_{0}^{2})}{1-c_{0}^{2}r_{0}^{2}}.$
So $B_{\rho}(c_{0},r_{0})\cap\mathbb{B}_{I}=B(c_{1},r_{1})\cap\mathbb{B}_{I}=\{x+Iy\in\mathbb{B}:|x+Iy-c_{1}|<r_{1}\}$ for all $I\in\mathbb{S}$. ∎
Proposition 3.5 (Quaternionic Dieudonné’s estimate 1).
Let $f:\mathbb{B}\longrightarrow\mathbb{B}$ be slice regular, such that $f(0)=0$. Suppose $q_{0}\in\mathbb{B}$ and set $\alpha_{f}(q_{0}):=\frac{1-|f(q_{0})|^{2}}{1-|q_{0}|^{2}}$. Then
(15) $\left|\alpha_{f}(q_{0})f^{h}(q_{0})-q_{0}^{-1}f(q_{0})\right|\leq\dfrac{|q_{0}|^{2}-|f(q_{0})|^{2}}{|q_{0}|(1-|q_{0}|^{2})}.$
Proof.
By (11) applied to the function $f^{\star}_{q_{0}}$, we have $\rho_{\mathbb{B}}(f^{h}(q_{0}),f^{\star}_{q_{0}}(0))\leq\rho_{\mathbb{B}}(q_{0},0)=|q_{0}|$: so by Lemma 3.1 $f^{h}(q_{0})$ lies in the pseudo-hyperbolic ball $B_{\rho}(q_{0}^{-1}f(q_{0}),|q_{0}|)$. By Lemma 3.4
$B_{\rho}(c_{0},r_{0})=B\left(c_{0}\frac{1-r_{0}^{2}}{1-|c_{0}|^{2}r_{0}^{2}},r_{0}\frac{1-|c_{0}|^{2}}{1-|c_{0}|^{2}r_{0}^{2}}\right),$
one gets that $f^{h}(q_{0})\in B\left(\alpha_{f}(q_{0})^{-1}q_{0}^{-1}f(q_{0}),\frac{|q_{0}|^{2}-|f(q_{0})|^{2}}{|q_{0}|(1-|f(q_{0})|^{2})}\right)$, which is equivalent to (15).
Corollary 3.6 (Quaternionic Dieudonné’s estimate 2).
By the same assumptions of Proposition 3.5,
(16) $\left|f^{h}(q_{0})\right|\leq\begin{cases}\alpha_{f}(q_{0})^{-1}&\text{if}\quad|q_{0}|\leq\sqrt{2}-1\\ \alpha_{f}(q_{0})^{-1}\dfrac{(1+|q_{0}|^{2})^{2}}{4|q_{0}|(1-|q_{0}|^{2})}&\text{if}\quad\sqrt{2}-1<|q_{0}|<1.\end{cases}$
Proof.
Inequalities (16) can be obtained from (15) by using exactly the same argument shown in [15, pag. 199-200]. ∎
Proposition 3.7 (Quaternionic Goluzin’s estimate).
Let $f:\mathbb{B}\longrightarrow\mathbb{B}$ be a slice regular self-map of the unit ball such that $f(0)=0$. Then for any $q_{0}\in\mathbb{B}$ it is
(17) $\left|f^{h}(q_{0})\right|\leq\dfrac{\left|\partial_{C}f(0)\right|+\frac{2|q_{0}|}{1+|q_{0}|^{2}}}{1+\left|\partial_{C}f(0)\right|\frac{2|q_{0}|}{1+|q_{0}|^{2}}}.$
Proof.
First we observe that
$f^{h}(0)=f^{\star}_{0}(0)=(\mathcal{M}_{0}^{-\ast}\ast(\mathcal{M}_{0}\bullet f))(0)=q^{-1}f(q)\mid_{q=0}=\partial_{C}f(0).$
We do our computation with the Poincaré distance: by the triangle inequality and (12) we have
$\delta_{\mathbb{B}}(0,f^{h}(q_{0}))\leq\delta_{\mathbb{B}}(0,f_{q_{0}}^{\star}(0))+\delta_{\mathbb{B}}(q_{0},0).$
Now the first term in the right hand side coincides with $\delta_{\mathbb{B}}(0,f_{0}^{\star}(q_{0}))$ by Lemma 3.1. By the same argument as above
$\delta_{\mathbb{B}}(0,f^{\star}_{0}(q_{0}))\leq\delta_{\mathbb{B}}(0,f^{h}(0))+\delta_{\mathbb{B}}(q_{0},0).$
In the end we get $\delta_{\mathbb{B}}(0,f^{h}(q_{0}))-\delta_{\mathbb{B}}(0,f^{h}(0))\leq 2\delta_{\mathbb{B}}(q_{0},0)$. In all the computations made above we can exchange 00 and $q_{0}$, thus
$\left|\delta_{\mathbb{B}}(0,f^{h}(q_{0}))-\delta_{\mathbb{B}}(0,f^{h}(0))\right|\leq 2\delta_{\mathbb{B}}(q_{0},0)=2\tanh^{-1}|q_{0}|,$
where the left hand side equals $\delta_{\mathbb{B}}(|f^{h}(q_{0})|,|\partial_{C}f(0)|)$. Applying the function $\tanh$ to both sides and recalling the duplication formula $\tanh(2x)=\tfrac{2\tanh x}{1+\tanh^{2}x}$, we obtain that
$\dfrac{|f^{h}(q_{0})|-|\partial_{C}f(0)|}{1-|\partial_{C}f(0)||f^{h}(q_{0})|}$ $\leq\rho_{\mathbb{B}}(|f^{h}(q_{0})|,|\partial_{C}f(0)|)\leq\dfrac{2|q_{0}|}{1+|q_{0}|^{2}}$
$\implies|f^{h}(q_{0})|-|\partial_{C}f(0)|$ $\leq\dfrac{2|q_{0}|}{1+|q_{0}|^{2}}-\left(|\partial_{C}f(0)|\dfrac{2|q_{0}|}{1+|q_{0}|^{2}}\right)|f^{h}(q_{0})|.$
By solving the last inequality in $|f^{h}(q_{0})|$ one gets (17).∎
3.1. Distortion estimates for the class $\boldsymbol{\mathcal{B}_{\alpha}}$
Let $\alpha\in[0,1)$ and let $\mathcal{B}_{\alpha}$ be the class of slice regular functions $f:\mathbb{B}\longrightarrow\mathbb{B}$ such that $f(0)=0$ and $\partial_{C}f(0)=\alpha$. Let us observe that in this case the Cullen derivative at zero coincides with the hyperbolic derivative $f^{h}(0)$ by the definition of the latter. For such functions we have the following estimates.
Proposition 3.8 (see [7, Thm 6.1]).
Let $10\leq\alpha<$ 1 and $f\in\mathcal{B}_{\alpha}$. Then for all $q\in\mathbb{B}$
$\dfrac{|f^{h}(q)-\alpha|}{|1-\alpha f^{h}(q)|}\leq\dfrac{2|q|}{1+|q|^{2}}$
and moreover
(18) $\dfrac{\alpha|q|^{2}-2|q|+\alpha}{|q|^{2}-2\alpha|q|+1}\leq\operatorname{Re}f^{h}(q)\leq|f^{h}(q)|\leq\dfrac{\alpha|q|^{2}+2|q|+\alpha}{|q|^{2}+2\alpha|q|+1}.$
The same estimates hold if we replace $f^{h}(q)$ with $f^{\star}_{q}(\overline{q})$.
Proof.
If $f\in\mathcal{B}_{\alpha}$, then it is not a Möbius transformation: otherwise $f(0)=0$ would imply that $f(q)=q\cdot u$, $u\in\partial\mathbb{B}$ and then it would be $\alpha=\partial_{C}f(0)=u$, a contradiction.
By applying the triangle inequality, recalling that by Lemma 3.1 $f^{\star}_{0}(q)=f^{\star}_{q}(0)$ and then using (12) twice one gets that
$\delta_{\mathbb{B}}(\alpha,f^{h}(q))$ $=\delta_{\mathbb{B}}(f^{\star}_{0}(0),f^{\star}_{q}(q))=\delta_{\mathbb{B}}(f^{\star}_{0}(0),f^{\star}_{0}(q))+\delta_{\mathbb{B}}(f^{\star}_{0}(q),f^{\star}_{q}(q))$
$=\delta_{\mathbb{B}}(f^{\star}_{0}(0),f^{\star}_{0}(q))+\delta_{\mathbb{B}}(f^{\star}_{q}(0),f^{\star}_{q}(q))\leq 2\delta_{\mathbb{B}}(0,q),$
which can be rewritten as
$\tanh^{-1}\rho_{\mathbb{B}}(\alpha,f^{h}(q))\leq 2\tanh^{-1}\rho_{\mathbb{B}}(0,q)=2\tanh^{-1}|q|;$
henceforth, by monotonicity of $\tanh$ and its duplication formula we have that
(19) $\rho_{\mathbb{B}}(\alpha,f^{h}(q))\leq\frac{2|q|}{1+|q|^{2}}$
and the left hand side corresponds by definition to $|M_{\alpha}(f^{h}(q))|=|1-\alpha f^{h}(q)|^{-1}|f^{h}(q)-\alpha|$.
Now, by Lemma 3.4 the inequality (19) says that for any $q\in\mathbb{B}$, $f^{h}(q)$ lies on a Euclidean ball contained in $\mathbb{B}$ which is symmetric with respect to $\mathbb{R}$ since $\alpha$ is real: the intersections with the real line of such ball are given by (14) where $c_{0}=\alpha$ and $r_{0}=2|q|/(1+|q|^{2})$:
$x_{\pm}=\dfrac{\alpha\pm 2|q|/(1+|q|^{2})}{1\pm 2\alpha|q|/(1+|q|^{2})}=\dfrac{\alpha|q|^{2}\pm 2|q|+\alpha}{|q|^{2}\pm 2\alpha|q|+1},\qquad x_{-}<x_{+},$
and its (Euclidean) center is
$\dfrac{x_{+}+x_{-}}{2}=\left[\dfrac{1-\left(\frac{2|q|}{1+|q|^{2}}\right)^{2}}{1-\alpha^{2}\left(\frac{2|q|}{1+|q|^{2}}\right)^{2}}\right]\cdot\alpha=\left[\dfrac{(1-|q|^{2})^{2}}{1+2(1-\alpha^{2})|q|^{2}+|q|^{4}}\right]\cdot\alpha\geq 0.$
Hence the real part of $f^{h}(q)$ lies between $x_{-}$ and $x_{+}$. Since $x_{+}\geq\max\{0,-x_{-}\}$, $x_{+}$ gives an upper bound also for $|f^{h}(q)|$.
If now we consider $f^{\star}_{q}(\overline{q})$, we have
$\delta_{\mathbb{B}}(\alpha,f^{\star}_{q}(\overline{q}))$ $=\delta_{\mathbb{B}}(f^{\star}_{0}(0),f^{\star}_{q}(\overline{q}))=\delta_{\mathbb{B}}(f^{\star}_{0}(0),f^{\star}_{0}(q))+\delta_{\mathbb{B}}(f^{\star}_{0}(q),f^{\star}_{q}(\overline{q}))$
$=\delta_{\mathbb{B}}(f^{\star}_{0}(0),f^{\star}_{0}(q))+\delta_{\mathbb{B}}(f^{\star}_{q}(0),f^{\star}_{q}(\overline{q}))\leq\delta_{\mathbb{B}}(0,q)+\delta_{\mathbb{B}}(0,\overline{q})=2\delta_{\mathbb{B}}(0,q)$
and so we get the same estimates by an identical argument. ∎
We recall that in the complex setting the estimate above can be used to determine the radius of injectivity of the analogous class $\mathcal{B}_{\alpha}$ ([7, Cor. 6.2]), given by $\varrho=\frac{\alpha}{1+\sqrt{1-\alpha^{2}}}$. The argument does not work anymore in our situation, since we can still deduce that for $|q|<\varrho$ it is $\operatorname{Re}f^{h}(q)>0$ and $\operatorname{Re}f^{\star}_{q}(\overline{q})>0$, hence by their expressions in terms of the Cullen and the spherical derivative of $f$ at $q$ – respectively given by (7) and (13) with $q$ in place of $p$ – it follows that both these derivatives are not vanishing. Differently from the complex case, this fact does not imply that the differential of the map $f$ is of maximal rank. So we should have got something more if we want to use the inverse function theorem to deduce (local) injectivity. However, $\frac{\alpha}{1+\sqrt{1-\alpha^{2}}}$ is also the injectivity radius of the quaternionic $\mathcal{B}_{\alpha}$. This is shown in [12] by a different approach.
4. Iterated hyperbolic difference quotients and the multipoint Schwarz–Pick Lemma
We have seen that if $f:\mathbb{B}\longrightarrow\mathbb{B}$ is slice regular and not a Möbius transformation, then for any $p_{1}\in\mathbb{B}$ the function $f^{\star}_{p_{1}}$ is a slice regular self-map of the unit ball. We can define the hyperbolic difference quotient of $f^{\star}_{p_{1}}$ with respect to $p_{2}\in\mathbb{B}$, getting a unimodular constant whether $f^{\star}_{p_{1}}$ is Möbius – hence $f$ is a regular Blaschke product of degree two – or obtaining a slice regular self-map of $\mathbb{B}$ otherwise. In the latter case we can go further with the iteration of such procedure taking the hyperbolic difference quotient with respect to a point $p_{3}$. In the end we have the following definition (see [13] or [6] for the complex case).
Definition 4.1.
Let $f:\mathbb{B}\longrightarrow\mathbb{B}$ be a slice regular function, let $n\in\mathbb{N},n>0$ and $p_{1},\ldots,p_{n}$ points in $\mathbb{B}$. Then we set $f^{\{1\}}_{p_{1}}(q):=f^{\star}_{p_{1}}(q)$ and for $n>1$
$f^{\{n\}}_{p_{1},\ldots,p_{n}}(q):=\begin{cases}f^{\{n-1\}}_{p_{1},\ldots,p_{n-1}}\quad\text{if }f^{\{n-1\}}_{p_{1},\ldots,p_{n-1}}\text{ is a unimodular constant,}\\ (\mathcal{M}_{p_{n}}^{-\ast}\ast(\mathcal{M}_{f^{\{n-1\}}_{p_{1},\ldots,p_{n-1}}(p_{n})}\bullet f^{\{n-1\}}_{p_{1},\ldots,p_{n-1}}))(q)\quad\text{ otherwise.}\end{cases}$
Let us observe that $f^{\{n\}}_{p_{1},\ldots,p_{n}}=(f^{\{n-1\}}_{p_{1},\ldots,p_{n-1}})^{\star}_{p_{n}}$ in the second case; moreover we get a unimodular constant in the case $f$ is a regular Blaschke product of degree less or equal to $n$, a Möbius transformation if $f$ is Blaschke of degree $n+1$ and a regular Blaschke product of degree $\kappa\geq 2$ if $f$ is Blaschke of degree $n+\kappa$.
Theorem 4.2.
Let $f:\mathbb{B}\longrightarrow\mathbb{B}$ be a slice regular function, $n>0$ integer and $p_{1},\ldots,p_{n}\in\mathbb{B}$. Then
$|f^{\{n\}}_{p_{1},\ldots,p_{n}}(q)|\leq 1\qquad\text{for all }q\in\mathbb{B}$
with strict inequality holding for each point, unless $f$ is a regular Blaschke product of degree $\leq n$, as the left term is a unimodular constant independent of $q$.
Proof.
It is Proposition 2.3 when $n=1$. For $n>1$, just apply the same result to the function $f^{\{n-1\}}_{p_{1},\ldots,p_{n-1}}$. ∎
Theorem 4.3 (Multipoint Schwarz–Pick Lemma).
Let $f:\mathbb{B}\longrightarrow\mathbb{B}$ be a slice regular function, $n>0$ integer and $p,p_{1},\ldots,p_{n}\in\mathbb{B}$. If $f$ is not a regular Blaschke product of degree $\leq n$, then
$|(\mathcal{M}_{f^{\{n\}}_{p_{1},\ldots,p_{n}}(p)}\bullet f^{\{n\}}_{p_{1},\ldots,p_{n}})(q)|\leq|\mathcal{M}_{p}(q)|$
for all $q\in\mathbb{B}$ and the inequality is strict for $q\in\mathbb{B}\setminus\{p\}$ unless $f$ is a regular Blaschke product of degree $n+1$.
Proof.
By Theorem 2.4 applied to $f^{\{n\}}_{p_{1},\ldots,p_{n}}$, which is a slice regular self-map of $\mathbb{B}$ by assumption. ∎
We can rewrite this result in terms of pseudo-hyperbolic metric:
$\rho_{\mathbb{B}}\left(f^{\{n\}}_{p_{1},\ldots,p_{n}}(p),f^{\{n\}}_{p_{1},\ldots,p_{n}}\left(\widetilde{\mathcal{T}}_{(f^{\{n\}}_{p_{1},\ldots,p_{n}}),p}(q)\right)\right)\leq\rho_{\mathbb{B}}(p,\mathcal{T}_{p}(q)),$
where $\mathcal{T}_{p}(q)=(1-qp)^{-1}q(1-qp)$ and
$\widetilde{\mathcal{T}}_{(f^{\{n\}}_{p_{1},\ldots,p_{n}}),p}(q):=h(q)^{-1}qh(q),\quad h(q)=1-(f^{\{n\}}_{p_{1},\ldots,p_{n}}(p)\ast(f^{\{n\}}_{p_{1},\ldots,p_{n}})^{c})(q).$
Let us observe that in the case $q\in\mathbb{R}\cap\mathbb{B}=(-1,1)$ the maps $\mathcal{T}_{p}$ and $\widetilde{\mathcal{T}}_{(f^{\{n\}}_{p_{1},\ldots,p_{n}}),p}$ act as identity on $q$ and we reduce to the classical Schwarz–Pick inequality in the pullback form:
$\rho_{\mathbb{B}}\left(f^{\{n\}}_{p_{1},\ldots,p_{n}}(p),f^{\{n\}}_{p_{1},\ldots,p_{n}}(q)\right)\leq\rho_{\mathbb{B}}(p,q).$
5. On Nevanlinna–Pick interpolation problem
A result on Nevanlinna–Pick interpolation problem over the unit ball of $\mathbb{B}$ which is essentially analogous to the complex case has already been stated by Alpay, Bolotnikov, Colombo and Sabadini in [2].
Theorem 5.1 ([2, Thm. 1.3]).
Given $n$ distinct points $p_{1},\ldots,p_{n}\in\mathbb{B}$ and $n$ values $s_{1},\ldots,s_{n}\in{\mathbb{B}}$, there exists a function $f\in\mathcal{SR}(\mathbb{B},{\mathbb{B}})$ such that $f(p_{m})=s_{m}$ for all $m=1,\ldots,n$ if and only if the Hermitian matrix
$P=\begin{pmatrix}\sum\limits_{\kappa=0}^{+\infty}p_{m}^{\kappa}(1-s_{m}\overline{s_{\ell}})\overline{p_{\ell}}^{\kappa}\end{pmatrix}_{m,\ell=1}^{n},$
namely the Pick matrix associated to $\{p_{i}\}$ and $\{s_{i}\}$, is positive semidefinite.
5.1. Two points Nevanlinna–Pick interpolation
When $n=2$, in the complex case the Pick matrix is positive semidefinite if and only if the ‘Schwarz–Pick’ condition holds:
$\left|\dfrac{s_{1}-s_{2}}{1-s_{1}\overline{s_{2}}}\right|\leq\left|\dfrac{p_{1}-p_{2}}{1-p_{1}\overline{p_{2}}}\right|.$
In the quaternionic framework we cannot expect an analogous result in general by the different expression of the Schwarz–Pick Lemma. However, if we assume that one of the point $p_{1},p_{2}$ is real, we can recover the same result.
Proposition 5.2.
Let $r\in(-1,1)$, $p\in\mathbb{B}\setminus\{r\}$ and $s,q\in{\mathbb{B}}$. There exists a slice regular function $f\in\mathcal{SR}(\mathbb{B},\mathbb{B})$ such that $f(r)=s$ and $f(p)=q$ if and only if
(20) $\rho_{\mathbb{B}}(q,s)\leq\rho_{\mathbb{B}}(p,r).$
Proof.
If there exists a function $f$ as required, then by (11) we recall that
$\rho_{\mathbb{B}}(f(p),f(\widetilde{\mathcal{T}}_{f,p}(r)))\leq\rho_{\mathbb{B}}(p,\mathcal{T}_{p}(r)),$
but $\widetilde{\mathcal{T}}_{f,p}(r)=\mathcal{T}_{p}(r)=r$ since $r$ is real: thus $\rho_{\mathbb{B}}(q,s)=\rho_{\mathbb{B}}(f(p),f(r))\leq\rho_{\mathbb{B}}(p,r)$.
Vice versa, let us assume $\rho_{\mathbb{B}}(q,s)\leq\rho_{\mathbb{B}}(p,r)$. Then, by noticing that
$1-\rho_{\mathbb{B}}(q,s)^{2}=1-|1-s\overline{q}|^{-2}|s-q|^{2}=|1-s\overline{q}|^{-2}(1+|s|^{2}|q|^{2}-|s|^{2}-|q|^{2})=\dfrac{(1-|s|^{2})(1-|q|^{2})}{|1-s\overline{q}|^{2}},$
our assumption is equivalent to say that
(21) $\dfrac{(1-|s|^{2})(1-|q|^{2})}{|1-s\overline{q}|^{2}}\geq\dfrac{(1-|r|^{2})(1-|p|^{2})}{|1-r\overline{p}|^{2}}.$
If we take the two pairs $(r,p)$ and $(s,q)$, the associated Pick matrix $P$ is such that
$P_{1,1}$ $=\sum\limits_{\kappa=0}^{+\infty}(1-|s|^{2})r^{\kappa}\overline{r}^{\kappa}=(1-|s|^{2})(1-|r|^{2})^{-1}$
$P_{2,2}$ $=\sum\limits_{\kappa=0}^{+\infty}(1-|q|^{2})p^{\kappa}\overline{p}^{\kappa}=(1-|q|^{2})(1-|p|^{2})^{-1}$
$P_{1,2}$ $=\sum\limits_{\kappa=0}^{+\infty}(1-s\overline{q})r^{\kappa}\overline{p}^{\kappa}=(1-s\overline{q})(1-r\overline{p})^{-1}$
$P_{2,1}$ $=\overline{P_{1,2}}=(1-pr)^{-1}(1-q\overline{s}).$
Clearly $P_{1,1}\geq 0$, so the matrix is positive semidefinite if and only if $P_{1,1}P_{2,2}-|P_{1,2}|^{2}\geq 0$, which is equivalent to (21). ∎
Now let’s assume that both $r$ and $p$ are real and $P\geq 0$. In this particular setting we can reproduce the argument of the complex case (see [6] for instance) to find the solutions of the interpolation problem.
Proposition 5.3.
Let $r,p\in(-1,1)$, $r\neq p$ and $s,q\in{\mathbb{B}}$, such that the Pick matrix associated to the pairs $(r,p)$ and $(s,q)$ is positive semidefinite, or equivalently (20) holds. If the Pick matrix is singular, then $Q:=M_{r}(p)^{-1}M_{s}(q)$ is unimodular and the interpolation problem
$f(r)=s,\quad f(p)=q,\qquad f\in\mathcal{SR}(\mathbb{B},\mathbb{B})$
has a unique solution given by the Möbius transformation
$\boxed{f=\mathcal{M}_{-s}\bullet(\mathcal{M}_{r}\ast Q).}$
Otherwise, if $P>0$, then $|Q|<1$ and the interpolation problem has infinite solutions parameterized by the space $\mathcal{SR}(\mathbb{B},\overline{\mathbb{B}})=\mathcal{SR}(\mathbb{B},\mathbb{B})\cup\partial\mathbb{B}$  [^4] This follows from the Maximum Modulus Principle in the quaternionic setting, see [16, Thm. 7.1].:
$\boxed{f=\mathcal{M}_{-s}\bullet(\mathcal{M}_{r}\ast(\mathcal{M}_{-Q}\bullet(\mathcal{M}_{p}\ast h)))}\qquad(h\in\mathcal{SR}(\mathbb{B},\mathbb{B})\cup\partial\mathbb{B})$
Proof.
By hypothesis $P\geq 0$, hence by Theorem 5.1 there exists a solution in $\mathcal{SR}(\mathbb{B},\mathbb{B})$ for the interpolation problem $f(r)=s$ and $f(p)=q$. But since we assumed that $r$ and $p$ are real, then $f^{\star}_{r}(p)=(\mathcal{M}_{r}^{-\ast}\ast(\mathcal{M}_{f(r)}\bullet f))(p)$ can be computed as $M_{r}(p)^{-1}M_{f(r)}(f(p))=M_{r}(p)^{-1}M_{s}(q)=Q$.
Consider first the singular case, $P\geq 0$ and $\det P=0$: in that case $\rho_{\mathbb{B}}(s,q)=\rho_{\mathbb{B}}(r,p)$, so $f^{\star}_{r}(p)=Q\in\partial\mathbb{B}$, thus $f^{\star}_{r}\equiv Q$ and $f^{\{2\}}_{r,p}\equiv Q$; therefore $f$ must be a Möbius transformation for which $\mathcal{M}_{r}^{-\ast}\ast(\mathcal{M}_{s}\bullet f)=Q$. It is easy to check that $f=\mathcal{M}_{-s}\bullet(\mathcal{M}_{r}\ast Q)$ is the solution, as $f(r)=M_{-s}(0)=s$ and $f(p)=M_{-s}(M_{r}(p)Q)=M_{-s}(M_{s}(q))=q$.
In the non-singular case $P>0$, we have that $g:=f^{\star}_{r}:\mathbb{B}\longrightarrow\mathbb{B}$ and sends $p$ to $Q\in\mathbb{B}$. Then $f^{\{2\}}_{r,p}=g^{\star}_{p}$ could be either a unimodular constant or a slice regular self-map of $\mathbb{B}$. It is straightforward to see that the problem
$\begin{cases}g^{\star}_{p}=\mathcal{M}_{p}^{-\ast}\ast(\mathcal{M}_{g(p)}\bullet g)=h\\ g(p)=Q\end{cases}$
has the unique slice regular solution $g=\mathcal{M}_{-Q}\bullet(\mathcal{M}_{p}\ast h)$ for any $h\in\mathcal{SR}(\mathbb{B},\mathbb{B})\cup\partial\mathbb{B}$; once we fix such a function $h$, we can find $f$ so that
$\begin{cases}f^{\star}_{r}=g=\mathcal{M}_{-Q}\bullet(\mathcal{M}_{p}\ast h)\\ f(r)=s\end{cases}$
in the same way, getting $f=\mathcal{M}_{-s}\bullet(\mathcal{M}_{r}\ast g)=\mathcal{M}_{-s}\bullet(\mathcal{M}_{r}\ast(\mathcal{M}_{-Q}\bullet(\mathcal{M}_{p}\ast h)))$. By direct computation, $f(p)=M_{-s}(M_{r}(p)g(p))=M_{-s}(M_{r}(p)Q)=M_{-s}(M_{s}(q))=q$. ∎
Example 5.4.
Let $\lambda,\mu\in(-1,1)$. We want to find conditions on these parameters such that there exists a slice regular map $f$ from $\mathbb{B}$ to itself for which $f\left(-\frac{1}{2}\right)=\lambda i$ and $f\left(\frac{1}{2}\right)=\mu j$. Then it is enough to compute
$Q=M_{-\frac{1}{2}}\left(\frac{1}{2}\right)^{-1}\cdot M_{\lambda i}(\mu j)=\frac{5}{4}(1-\lambda\mu k)^{-1}(\mu j-\lambda i)\implies|Q|^{2}=\frac{25}{16}\cdot\frac{\mu^{2}+\lambda^{2}}{1+\lambda^{2}\mu^{2}}.$
So the condition for the existence of a solution for the interpolation problem $|Q|\leq 1$ in this case becomes $(25-16\lambda^{2})\mu^{2}\leq 16-25\lambda^{2}$. In particular, if we fix $\lambda$ so that $1>|\lambda|>\frac{4}{5}$, we have no solutions for any choice of $\mu$; if otherwise $\lambda$ is fixed so that $0\leq|\lambda|\leq\frac{4}{5}$, then we have solutions once
$|\mu|\leq\sqrt{\frac{(4+5\lambda)(4-5\lambda)}{(5+4\lambda)(5-4\lambda)}}.$
When equality holds ($|Q|=1$), we have the extremal case of a unique (Möbius) solution, otherwise we have infinite solutions parameterized by $\mathcal{SR}(\mathbb{B},{\mathbb{B}})\cup\partial\mathbb{B}$.
For instance, let $\lambda=\frac{1}{4}$. Then we have solutions if and only if $|\mu|\leq\sqrt{\frac{77}{128}}$. When the inequality is strict, solutions are given (implicitly) by the parametric formula
$f=\mathcal{M}_{-\frac{1}{4}i}\bullet(\mathcal{M}_{-\frac{1}{2}}\ast(\mathcal{M}_{-Q}\bullet(\mathcal{M}_{\frac{1}{2}}\ast h))),\quad Q=\frac{5}{4}(4-\mu k)^{-1}(-i+4\mu j),\quad h\in\mathcal{SR}(\mathbb{B},{\mathbb{B}})\cup\partial\mathbb{B}.$
When $\mu=\pm\sqrt{\frac{77}{128}}$, the unique solution is $f=\mathcal{M}_{-\frac{1}{4}i}\bullet(\mathcal{M}_{-\frac{1}{2}}\ast Q)$.
5.2. Nevanlinna–Pick interpolation with three points
For the case $n\geq 3$, in the complex setting, a result which was stated by Beardon and Minda [7, Thm. 7.1] in the case $n=3$ and then generalized by Baribeau, Rivard and Wegert [6, Thm. 4.1–4.2] gives a more geometric set of conditions which are equivalent to the existence of solutions for the Nevanlinna–Pick interpolation problem. If we assume that the nodes in the domain are real, we get exactly the same conditions also in our framework.
For the sake of simplicity, we first see the case $n=3$ before going to the general picture.
Theorem 5.5.
Let $r_{1},r_{2},r_{3}\in(-1,1)$ be distinct and let $s_{1},s_{2},s_{3}\in\mathbb{B}$. Consider the interpolation problem
$f(r_{m})=s_{m}\quad\text{for all }m=1,2,3,\qquad f\in\mathcal{SR(\mathbb{B},\mathbb{B})}.$
Let us define
$Q_{1}^{2}$ $:=M_{r_{1}}(r_{2})^{-1}M_{s_{1}}(s_{2})$
$Q_{1}^{3}$ $:=M_{r_{1}}(r_{3})^{-1}M_{s_{1}}(s_{3})$
and  [^5] Here $\infty$ is just a conventional definition, it does not correspond to a pole of the ratio. For our purposes, we could have chosen any quaternion of norm $>1$. The other cases correspond to having as $Q_{1}^{2},Q_{1}^{3}$ two different points on $\partial\mathbb{B}$, or a point inside $\mathbb{B}$ and the other one outside, or at least one of the two points not in $\overline{\mathbb{B}}$.
$Q_{2}^{3}:=\begin{cases}M_{r_{2}}(r_{3})^{-1}M_{Q_{1}^{2}}(Q_{1}^{3})&\text{ if }Q_{1}^{2},Q_{1}^{3}\in\mathbb{B},\\ Q_{1}^{2}&\text{ if }Q_{1}^{2}=Q_{1}^{3}\in\partial\mathbb{B},\\ \infty&\text{ otherwise}.\end{cases}$
  • If $|Q_{2}^{3}|<1$, we have an infinite family of solutions (non-singular case) given by
    (22) $\boxed{\mathcal{M}_{-s_{1}}\bullet(\mathcal{M}_{r_{1}}\ast(\mathcal{M}_{-Q_{1}^{2}}\bullet(\mathcal{M}_{r_{2}}\ast(\mathcal{M}_{-Q_{2}^{3}}\bullet(\mathcal{M}_{r_{3}}\ast h)))))}$
    where $h$ is a function in $\mathcal{SR}(\mathbb{B},{\mathbb{B}})$ or a unimodular constant. In the last situation, the solution is a regular Blaschke product of degree three.
  • If $Q_{2}^{3}\in\partial\mathbb{B}$, there are two cases.
    1. a)
      If $Q_{1}^{2},Q_{1}^{3}$ both belong to $\mathbb{B}$, there exists a unique solution (singular case) given by the regular Blaschke product of degree two
      (23) $\boxed{\mathcal{M}_{-s_{1}}\bullet(\mathcal{M}_{r_{1}}\ast(\mathcal{M}_{-Q_{1}^{2}}\bullet(\mathcal{M}_{r_{2}}\ast Q_{2}^{3})))}$
    2. b)
      When $Q_{1}^{2}=Q_{1}^{3}\in\partial\mathbb{B}$, the solution is given by the regular Möbius map
      (24) $\boxed{\mathcal{M}_{-s_{1}}\bullet(\mathcal{M}_{r_{1}}\ast Q_{1}^{2})}$
  • Otherwise, if $Q_{2}^{3}$ has modulus strictly greater than one, the interpolation problem has no solutions.
Proof.
We first observe that if we have a (slice regular) solution $f:\mathbb{B}\longrightarrow\mathbb{B}$ of the interpolation problem $f(r_{m})=s_{m},m=1,2,3$, then
(25) $Q_{1}^{2}=f^{\star}_{r_{1}}(r_{2}),\quad Q_{1}^{3}=f^{\star}_{r_{1}}(r_{3}),\quad Q_{2}^{3}=f^{\{2\}}_{r_{1},r_{2}}(r_{3}).$
Indeed $f^{\star}_{r_{1}}\in\mathcal{SR}(\mathbb{B},{\mathbb{B}})\cup\partial\mathbb{B}$, hence $\overline{\mathbb{B}}\ni f^{\star}_{r_{1}}(r_{m})=\mathcal{M}_{r_{1}}^{-\ast}\ast(\mathcal{M}_{s_{1}}\bullet f)\mid_{r_{m}}$, $m=2,3$: since $r_{m}$ is real, the last term is equal to $\mathcal{M}_{r_{1}}(r_{m})^{-1}(\mathcal{M}_{s_{1}}\bullet f)(r_{m})=M_{r_{1}}(r_{m})^{-1}M_{s_{1}}(f(r_{m}))=M_{r_{1}}(r_{m})^{-1}M_{s_{1}}(s_{m})=Q_{1}^{m}$ for $m=2,3$. If $f$ is a Möbius transformation, then $f^{\star}_{r_{1}}=f^{\{2\}}_{r_{1},r_{2}}\equiv\eta\in\partial\mathbb{B}$, so we have also $Q_{1}^{2}=Q_{1}^{3}=Q_{2}^{3}=\eta$. Otherwise, $f^{\star}_{r_{1}}$ has values in the open unit ball and $f^{\{2\}}_{r_{1},r_{2}}(r_{3})=(f^{\star}_{r_{1}})^{\star}_{r_{2}}(r_{3})$, which as above can be computed as $M_{r_{2}}^{-1}(r_{3})M_{f^{\star}_{r_{1}}(r_{2})}(f^{\star}_{r_{1}}(r_{3}))=M_{r_{2}}^{-1}(r_{3})M_{Q_{1}^{2}}(Q_{1}^{3})=Q_{2}^{3}$.
This claim shows that we cannot have solutions in the case $Q_{2}^{3}$ has modulus strictly greater than one.
Let us assume $|Q_{2}^{3}|<1$. This implies by its definition that also $Q_{1}^{2},Q_{1}^{3}$ lie within the open unit ball. Hence by (25), if $f$ exists, then both $f^{\star}_{r_{1}}$ and $f^{\{2\}}_{r_{1},r_{2}}$ are slice regular maps from $\mathbb{B}$ to itself. Moreover we can define $f^{\{3\}}_{r_{1},r_{2},r_{3}}\in\mathcal{SR(\mathbb{B},\mathbb{B})}\cup\partial\mathbb{B}$. Now, let’s take $h$ arbitrarily in this latter set of functions and seek for a function $g_{1}$ whose hyperbolic difference quotient at $r_{3}$ is $h$ and attains the value $Q_{2}^{3}$ at $r_{3}$, i.e., the solution of
$\begin{cases}(g_{1})^{\star}_{r_{3}}=h\\ g_{1}(r_{3})=Q_{2}^{3}\end{cases}.$
But $(g_{1})^{\star}_{r_{3}}=h\iff\mathcal{M}_{g_{1}(r_{3})}\bullet g_{1}=\mathcal{M}_{r_{3}}\ast h$, hence the problem above has a unique (slice regular) solution given by $g_{1}=\mathcal{M}_{-Q_{2}^{3}}\bullet(\mathcal{M}_{r_{3}}\ast h)$. Now we proceed backwards, finding $g_{2}$ such that
$\begin{cases}(g_{2})^{\star}_{r_{2}}=g_{1}\\ g_{2}(r_{2})=Q_{1}^{2}\end{cases}\implies g_{2}=\mathcal{M}_{-Q_{1}^{2}}\bullet(\mathcal{M}_{r_{2}}\ast g_{1}).$
By (25) we should also have $g_{2}(r_{3})=Q_{1}^{3}$, but evaluating the solution above at $r_{3}$ one directly gets $g_{2}(r_{3})=M_{-Q_{1}^{2}}(M_{r_{2}}(r_{3})\cdot Q_{2}^{3})=M_{-Q_{1}^{2}}(M_{Q_{1}^{2}}(Q_{1}^{3}))=Q_{1}^{3}$. It remains to find $g_{3}$ such that
$\begin{cases}(g_{3})^{\star}_{r_{1}}=g_{2}\\ g_{3}(r_{1})=s_{1}\end{cases}\implies g_{3}=\mathcal{M}_{-s_{1}}\bullet(\mathcal{M}_{r_{1}}\ast g_{2}).$
In the end, we get a function $f=g_{3}$, depending on the choice of $h$, which is slice regular from $\mathbb{B}$ to itself and satisfies $f(r_{1})=s_{1}$; moreover, by direct evaluation, for $m=2,3$ it is $f(r_{m})=M_{-s_{1}}(M_{r_{1}}(r_{m})\cdot g_{2}(r_{m}))=M_{-s_{1}}(M_{r_{1}}(r_{m})\cdot Q_{1}^{m})=M_{-s_{1}}(M_{s_{1}}(s_{m}))=s_{m}$. Henceforth, it satisfies the interpolation conditions.
We can summarize this algorithm by the following scheme (we indicate with bold symbols the conditions we impose at each step, the others follow automatically):
$r_{1}$ $r_{2}$ $r_{3}$
$\boldsymbol{s_{1}}$ $s_{2}$ $s_{3}$ $f=g_{3}$
$\boldsymbol{Q_{1}^{2}}$ $Q_{1}^{3}$ $f^{\star}_{r_{1}}=g_{2}$
$\boldsymbol{Q_{2}^{3}}$ $f^{\{2\}}_{r_{1},r_{2}}=g_{1}$
$f^{\{3\}}_{r_{1},r_{2},r_{3}}=h$
Now suppose $|Q_{2}^{3}|=1$, i.e., it is a point in the unit sphere $\partial\mathbb{B}$. There are two cases:
  1. a)
    $Q_{1}^{2}$, $Q_{1}^{3}$ both lie in $\mathbb{B}$: if $f$ exists, by (25) it is $f^{\star}_{r_{1}}\in\mathcal{SR}(\mathbb{B},\mathbb{B})$, while $f^{\{2\}}_{r_{1},r_{2}}$ is constant and equal to $Q_{2}^{3}$. So this time we start by finding the slice regular function whose hyperbolic difference quotient at $r_{2}$ is $Q_{2}^{3}$ and assumes the value $Q_{1}^{2}$ at $r_{2}$, namely $\widetilde{g}=\mathcal{M}_{-Q_{1}^{2}}\bullet(\mathcal{M}_{r_{2}}\ast Q_{2}^{3})$: by direct evaluation and since in our case $Q_{2}^{3}=M_{r_{2}}(r_{3})$ it also satisfies $\widetilde{g}(r_{3})=Q_{1}^{3}$. Since it is the result of a Möbius transformation acting on another Möbius transformation by $\bullet$, the map $\widetilde{g}$ is Möbius too. Finally we find $f$ for which $f^{\star}_{r_{1}}=\widetilde{g}$ and $f(r_{1})=s_{1}$, that is to say the second-degree regular Blaschke product $f=\mathcal{M}_{-s_{1}}\bullet(\mathcal{M}_{r_{1}}\ast\widetilde{g})$, which also satisfies $f(r_{2})=s_{2}$ and $f(r_{3})=s_{3}$.
  2. b)
    $Q_{1}^{2}=Q_{1}^{3}\in\partial\mathbb{B}$. In this latter case, only the last step of the algorithm is needed, since $f^{\star}_{r_{1}}$ should be the unimodular constant $u=Q_{1}^{2}=Q_{1}^{3}$. Then $f=\mathcal{M}_{-s_{1}}\bullet(\mathcal{M}_{r_{1}}\ast u)$, which is now a Möbius transformation, does the job by the same arguments seen above.∎
Remark 5.6.
We observe that the terms $Q_{\kappa}^{\ell}$ of the algorithm depend on the order we fix for the nodes of interpolation $r_{1},r_{2},r_{3}$, but the problem itself should not depend on that order. Hence, although $Q_{2}^{3}$ could change, the property of being inside, outside or in the boundary of the unit ball $\mathbb{B}$ is preserved. Moreover, a permutation of nodes produces a different expression for the family of solutions (22) in the non-singular case and the unique solution (23), or (24), in the singular one: this is not a contradiction as those representations in terms of $\ast$-products and $\bullet$-actions are not unique.
Example 5.7.
Let $\lambda,\mu\in(-1,1)$. For which value of these parameters there exists a slice regular map $f$ from $\mathbb{B}$ to itself for which $f(0)=0$, $f\left(-\frac{1}{2}\right)=\lambda i$ and $f\left(\frac{1}{2}\right)=\mu j$?
As seen before, we can solve this problem by computing:
$Q_{1}^{2}$ $=M_{0}\left(-\frac{1}{2}\right)^{-1}\cdot M_{0}(\lambda i)=-2\lambda i$
$Q_{1}^{3}$ $=M_{0}\left(\frac{1}{2}\right)^{-1}\cdot M_{0}(\mu j)=2\mu j$
$Q_{2}^{3}$ =$=\begin{cases}M_{-\frac{1}{2}}\left(\frac{1}{2}\right)^{-1}\cdot M_{Q_{1}^{2}}(Q_{1}^{3})=\frac{5}{2}(1+4\lambda\mu k)^{-1}(\lambda i+\mu j)\quad\text{if }|\lambda|,|\mu|<\frac{1}{2}\\ \infty\quad\text{otherwise}.\end{cases}$
The case $Q_{2}^{3}=Q_{1}^{2}\in\partial\mathbb{B}$ never happens since $Q_{1}^{3}$ is different from $Q_{1}^{2}$ for any choice of $\lambda$ and $\mu$, unless $\lambda=\mu=0$, but then they are equal to 00 and they’re not in the boundary of $\mathbb{B}$.
Now $|Q_{2}^{3}|^{2}=\frac{25}{4}\cdot\frac{\lambda^{2}+\mu^{2}}{1+16\lambda^{2}\mu^{2}}$. So $|Q_{2}^{3}|\leq 1$ if and only if $(25-64\lambda^{2})\mu^{2}\leq 4-25\lambda^{2}$ for $\lambda,\mu\in(-\frac{1}{2},\frac{1}{2})$. Suppose $\lambda$ is fixed so that $\frac{1}{2}>|\lambda|>\frac{2}{5}$: then we have no solutions for any choice of $\mu$; if otherwise we fix $\lambda$ such that $0\leq|\lambda|\leq\frac{2}{5}$, then $025-64\lambda^{2}>$ 0 and we conclude that the interpolation problem has solutions for
$|\mu|\leq\sqrt{\frac{(2+5\lambda)(2-5\lambda)}{(5+8\lambda)(5-8\lambda)}}.$
The problem is singular when the inequalities above are equalities.
For instance, if $\lambda=1/4$, then we can find solutions once $|\mu|\leq\sqrt{\frac{13}{112}}=\frac{1}{4}\sqrt{\frac{13}{7}}$. If the strict inequality holds, there is a family of solutions of the form
$q\ast(\mathcal{M}_{\frac{i}{2}}\bullet(\mathcal{M}_{-\frac{1}{2}}\ast(\mathcal{M}_{-Q_{2}^{3}}\bullet(\mathcal{M}_{\frac{1}{2}}\ast h)))),\quad Q_{2}^{3}=\frac{5}{2}(1+\mu k)^{-1}\cdot\left(\frac{1}{4}i+\mu j\right),\quad h\in\mathcal{SR}(\mathbb{B},{\mathbb{B}})\cup\partial\mathbb{B},$
while in the singular cases $\mu=\pm\sqrt{\frac{13}{112}}$ the unique solution is the regular Blaschke product of degree two $q\ast(\mathcal{M}_{\frac{i}{2}}\bullet(\mathcal{M}_{-\frac{1}{2}}\ast Q_{2}^{3}))$.
Corollary 5.8.
Let $r_{1},r_{2},r_{3}$ be pairwise distinct points in $(-1,1)$ and let $s_{1},s_{2},s_{3}\in\mathbb{B}$. The interpolation problem $f(r_{m})=s_{m},m=1,2,3$ with $f\in\mathcal{SR}(\mathbb{B},\mathbb{B})$ has infinite solutions if and only if
(26) $\rho_{\mathbb{B}}(s_{1},s_{m})<\rho_{\mathbb{B}}(r_{1},r_{m})\qquad\text{for }m=2,3$
and
(27) $\rho_{\mathbb{B}}(M_{r_{1}}(r_{2})^{-1}M_{s_{1}}(s_{2}),M_{r_{1}}(r_{3})^{-1}M_{s_{1}}(s_{3}))<\rho_{\mathbb{B}}(r_{2},r_{3}).$
If equality holds in (27), we have a unique solution given by a regular Blaschke product of degree two.
Inequalities (26) imply that the two points in the left hand side of (27) both lie in the open unit ball, so it makes sense to consider their pseudo-hyperbolic distance in $\mathbb{B}$.
Proof.
By Theorem 5.5, the interpolation problem has infinite solutions if and only if $|Q_{2}^{3}|<1$, and that’s true if and only if $Q_{1}^{2},Q_{1}^{3}\in\mathbb{B}$ and $|M_{Q_{1}^{2}}(Q_{1}^{3})|<|M_{r_{2}}(r_{3})|$, which correspond to the strict inequalities (26) and (27).
If (26) holds and we have equality in place of strict inequality in (27), then this corresponds to $Q_{1}^{2},Q_{1}^{3}\in\mathbb{B}$ with $|Q_{2}^{3}|=1$ and we still conclude by Theorem 5.5. ∎
5.3. Nevanlinna–Pick interpolation with more than three points
The same procedure by iteration seen above could be used to deal with the case $n>3$, as illustrated in [6] for the complex case.
Theorem 5.9.
Let us assume $n\geq 3$, let $r_{1},r_{2},\ldots,r_{n}\in(-1,1)$ be distinct and let $s_{1},s_{2},\ldots s_{n}\in\mathbb{B}$. Consider the interpolation problem
(28) $f(r_{m})=s_{m}\quad\text{for all }m=1,\ldots,n,\qquad f\in\mathcal{SR(\mathbb{B},\mathbb{B})}.$
Let
$Q_{0}^{m}:=s_{m}\qquad\text{for all }m=1,\ldots,n$
$Q_{1}^{m}:=M_{r_{1}}(r_{m})^{-1}M_{s_{1}}(s_{m})\qquad\text{for all }m=2,\ldots,n$
and for any $\kappa=2,\ldots,n-1$, $\ell=\kappa+1,\ldots,n$
$Q_{\kappa}^{\ell}:=\begin{cases}M_{r_{\kappa}}(r_{\ell})^{-1}M_{Q_{\kappa-1}^{\kappa}}(Q_{\kappa-1}^{\ell})&\text{ if }Q_{\kappa-1}^{\kappa},Q_{\kappa-1}^{\ell}\in\mathbb{B},\\ Q_{\kappa-1}^{\kappa}&\text{ if }Q_{\kappa-1}^{\kappa}=Q_{\kappa-1}^{\ell}\in\partial\mathbb{B},\\ \infty&\text{ otherwise}.\end{cases}$
The interpolation problem is non-singular, with infinite solutions, when $|Q_{n-1}^{n}|<1$. It is singular, with a unique (regular Blaschke) solution, if $|Q_{n-1}^{n}|=1$, whose degree is $\kappa_{0}\in\{1,\ldots,n-1\}$ if and only if $Q_{\kappa_{0}}^{\kappa_{0}+1}=\ldots=Q_{\kappa_{0}}^{n}\in\partial\mathbb{B}$, but $Q_{\kappa_{0}-1}^{\ell}\in\mathbb{B}$ for some (hence any) $\ell\in\{\kappa_{0},\kappa_{0}+1,\ldots,n\}$. Otherwise, the problem has no admissible solutions.
Proof.
The proof is essentially the same as in the case of three points, just with more iterations of the algorithm, which we can illustrate by the following scheme.
$r_{1}$ $r_{2}$ $r_{3}$cdots $r_{n-1}$ $r_{n}$
$\boldsymbol{s_{1}}$ $s_{2}$ $s_{3}$cdots $s_{n-1}$ $s_{n}$ $f=g_{n}$
$\boldsymbol{Q_{1}^{2}}$ $Q_{1}^{3}$cdots $Q_{1}^{n-1}$ $Q_{1}^{n}$ $f^{\star}_{r_{1}}=g_{n-1}$
$\boldsymbol{Q_{2}^{3}}$cdots $Q_{2}^{n-1}$ $Q_{2}^{n}$ $f^{\{2\}}_{r_{1},r_{2}}=g_{n-2}$
$\boldsymbol{\ddots}$vdots $\vdots$ $\vdots$
$\boldsymbol{Q_{n-2}^{n-1}}$ $Q_{n-2}^{n}$ $f^{\{n-2\}}_{r_{1},\ldots,r_{n-2}}=g_{2}$
$\boldsymbol{Q_{n-1}^{n}}$ $f^{\{n-1\}}_{r_{1},\ldots,r_{n-1}}=g_{1}$
$f^{\{n\}}_{r_{1},\ldots,r_{n}}=h$
Analogously to the three-point case, the existence of an admissible interpolating function $f$ implies that
(29) $f^{\{\kappa\}}_{r_{1},\ldots,r_{\kappa}}(r_{\ell})=Q_{\kappa}^{\ell}\qquad\kappa\in\{1,\ldots,n-1\},\ell\in\{\kappa+1,\ldots,n\}.$
This forces in particular $|Q_{n-1}^{n}|\leq 1$ in order to have a solution.
Assuming $|Q_{n-1}^{n}|<1$, by the definition of the coefficients in the scheme, all of them must belong to the unit ball $\mathbb{B}$, and then (29) implies that for a slice regular interpolating function $f:\mathbb{B}\longrightarrow\mathbb{B}$ – if existing – we have $f^{\{n-1\}}_{r_{1},\ldots,r_{n-1}}:\mathbb{B}\longrightarrow\mathbb{B}$ slice regular and $f^{\{n\}}_{r_{1},\ldots,r_{n}}\in\mathcal{SR}(\mathbb{B},\mathbb{B})$. We proceed following the scheme from the bottom to the top: given $h\in\mathcal{SR}(\mathbb{B},{\mathbb{B}})$ or $h\equiv u\in\partial\mathbb{B}$, we find $g_{1}$ such that $(g_{1})^{\star}_{r_{n}}=h$ and $g_{1}(r_{n})=Q^{n}_{n-1}$, then $g_{2}$ for which $(g_{2})^{\star}_{r_{n-1}}=g_{1}$ and $g_{2}(r_{n-1})=Q_{n-2}^{n-1}$: the solution will also satisfy $g_{2}(r_{n})=Q_{n-2}^{n}$. Following the rows, in the end one gets a slice regular function $f=g_{n}$ satisfying (29) $f(r_{m})=s_{m}$ for each $m=1,\ldots,n$, depending on the choice of the parameter $h$.
Namely, we get
$\boxed{f=g_{n}=\mathcal{M}_{-s_{1}}\bullet(\mathcal{M}_{r_{1}}\ast(\overbrace{\mathcal{M}_{-Q_{1}^{2}}\bullet(\mathcal{M}_{r_{2}}\ast(\ldots\ast(\underbrace{\mathcal{M}_{-Q_{n-1}^{n}}\bullet(\mathcal{M}_{r_{n}}\ast h)}_{g_{1}})))}^{g_{n-1}}))}$
If now we assume $|Q_{n-1}^{n}|=1$, i.e, $Q_{n-1}^{n}=u\in\partial\mathbb{B}$ this means – as a consequence of the definition of the terms $Q_{\kappa}^{\ell}$ – that all $Q_{\kappa}^{\ell}=u$ for all indices $\kappa$ greater or equal some $\kappa_{0}\in\{1,\ldots,n-1\}$, i.e., all the terms from the $(\kappa_{0}+1)$-th row of the scheme to the bottom. Suppose that the $\kappa_{0}$-th row contains a term within the unit ball: since the terms $Q_{\kappa_{0}}^{\ell}$ are not $\infty$, all the $\kappa_{0}$-th row consists of points of $\mathbb{B}$. By (29), for an interpolating function $f$ we have that $f^{\star}_{r_{1}},f^{\{2\}}_{r_{1},r_{2}},\ldots f^{\{\kappa_{0}-1\}}_{r_{1},\ldots,r_{\kappa_{0}-1}}\in\mathcal{SR}(\mathbb{B},\mathbb{B})$ while $f^{\kappa_{0}}_{r_{1},\ldots,r_{\kappa_{0}}}\equiv u\in\partial\mathbb{B}$. So we start from that row and go upwards in the scheme: we find $g_{n-\kappa_{0}+1}$ such that $(g_{n-\kappa_{0}+1})^{\star}_{r_{\kappa_{0}}}=g_{n-\kappa_{0}}\equiv u$ and $g_{n-\kappa_{0}+1}(r_{\kappa_{0}})=Q_{\kappa_{0}-1}^{\kappa_{0}}$, i.e., $g_{n-\kappa_{0}+1}=\mathcal{M}_{-Q_{\kappa_{0}-1}^{\kappa_{0}}}\bullet(\mathcal{M}_{r_{\kappa_{0}}}\ast u)$ and so on, until we get
$\boxed{f=g_{n}=\mathcal{M}_{-s_{1}}\bullet(\mathcal{M}_{r_{1}}\ast(\overbrace{\mathcal{M}_{-Q_{1}^{2}}\bullet(\mathcal{M}_{r_{2}}\ast(\ldots\ast(\underbrace{\mathcal{M}_{-Q_{\kappa_{0}-1}^{\kappa_{0}}}\bullet(\mathcal{M}_{r_{\kappa_{0}}}\ast u)}_{g_{n-\kappa_{0}+1}})))}^{g_{n-1}})).}$
which is a regular Blaschke product of degree $\kappa_{0}$ by Proposition 1.22; moreover it satisfies (29) and the interpolation condition $f(r_{m})=s_{m}$, $m=1,\ldots,n$.
The condition for having an infinite family of solutions (non-singular case) for the Pick interpolation problem is $|Q_{n-1}^{n}|<1$, or equivalently $|Q_{\kappa}^{\ell}|<1$ for all $\kappa=1,\ldots,n-1$ and $\ell=\kappa+1,\ldots,n$. This can be easily restated in terms of the pseudo-hyperbolic distance $\rho_{\mathbb{B}}$, analogously as in Corollary 5.8.
Corollary 5.10.
The interpolation problem (28) has infinite solutions if and only if
$\rho_{\mathbb{B}}(Q_{n-2}^{n-1},Q_{n-2}^{n})<\rho_{\mathbb{B}}(r_{n-1},r_{n}),$
or equivalently, if $\rho_{\mathbb{B}}(Q_{\kappa-1}^{\kappa},Q_{\kappa-1}^{\ell})<\rho_{\mathbb{B}}(r_{\kappa},r_{\ell})$ for all $\kappa\in\{1,\ldots,n-1\},\ell\in\{\kappa+1,\ldots,n\}$ – where we use the notation $Q_{0}^{\ell}:=s_{\ell}$.
5.4. The case of nodes not in the real line
Since now we have considered a particular case for the quaternionic Nevanlinna–Pick interpolation, when all the nodes are real. One may ask if Theorem 5.9 still works by replacing $r_{1},\ldots,r_{n}\in(-1,1)$ with distinct points $p_{1},\ldots p_{n}$ in the whole unit ball.
In generality, the argument shown above breaks down. Indeed, we would like to define from the data of the interpolation problem $\{p_{m}\},\{s_{m}\}$ quantities $Q_{\kappa}^{\ell}$ such that $f^{\star}_{p_{1}}(p_{\ell})=Q_{1}^{\ell}$ ($\ell=2,\ldots,n$) for any solution $f$ of the problem. But, assuming that $f$ exists,
$f^{\star}_{p_{1}}(p_{\ell})=(\mathcal{M}_{p_{1}}^{-\ast}\ast\mathcal{M}_{s_{1}}\bullet f)(p_{\ell})$
and the last term evaluates as $M_{p_{1}}(\tilde{p}_{\ell})^{-1}M_{s_{1}}(f(\hat{p}_{\ell}))$ where $\tilde{p}_{\ell}$ is a point in $\mathbb{S}_{p_{\ell}}$ depending on $p_{1}$ and $\hat{p}_{\ell}\in\mathbb{S}_{p_{\ell}}$ may depend on $p_{1},s_{1}$ and on the conjugate function $f^{c}$. They trivially coincide with $p_{\ell}$ when that point is real. So, the proposed algorithm does not work anymore, as $\hat{p}_{\ell}$ generally depends on the function $f^{c}$ which is not known a priori and in any case $f(\hat{p}_{\ell})$ is not known if $\hat{p}_{\ell}$ is not one of the nodes.
However, there is a specific situation in which we can recover a partial result in the spirit of Theorem 5.9. As Theorem 5.5 holds in the complex setting ([6, Thm. 4.1]) without the restriction to real nodes of interpolation, then we have the following partial answer: if $p_{1},\ldots,p_{n}\in\mathbb{B}$ belong to a common slice $\mathbb{C}_{I}$ and also $s_{1},\ldots,s_{n}\in\mathbb{B}\cap\mathbb{C}_{I}$, then the interpolation problem $f(p_{m})=s_{m}$, $m\in\{1,\ldots,n\}$ admits infinite solutions that are $\mathbb{C}_{I}$-preserving if $|Q_{n-1}^{n}|<1$, while it admits at least a solution given by a regular Blaschke product of degree at most $n-1$, when $|Q_{n-1}^{n}|=1$, and this is the unique one preserving $\mathbb{C}_{I}$.
Indeed, we can identify $\mathbb{B}\cap\mathbb{C}_{I}$ with the complex unit disk $\mathbb{D}$. Notice that $Q_{n-1}^{n}\in\mathbb{C}_{I}$ by our assumptions. If $|Q_{n-1}^{n}|\leq 1$, there exists a holomorphic function $f_{0}:\mathbb{B}\cap\mathbb{C}_{I}\longrightarrow\mathbb{B}\cap\mathbb{C}_{I}$ satisfying the interpolation properties. Now there exists a unique slice regular function $f:\mathbb{B}\longrightarrow\mathbb{H}$ extending $f_{0}$ by the formula (see [16, Lemma 1.21]):
$f(x+yJ)=\dfrac{f_{0}(x-yI)+f_{0}(x+yI)}{2}+\dfrac{JI[f_{0}(x-yI)-f_{0}(x+yJ)]}{2}$
$\forall\,x,y\in\mathbb{R},x^{2}+y^{2}<1,J\in\mathbb{S}.$
Fix $J$ as above and let $\lambda=\operatorname{Re}(JI)$. Then $|\lambda|\leq 1$ and there exists $K$ orthogonal to $I$ such that $JI=\lambda+\sqrt{1-\lambda^{2}}K$, so
$f(x+yJ)$ $=\dfrac{(1+\lambda)f_{0}(x-yI)+(1-\lambda)f_{0}(x+yI)}{2}+K\dfrac{\sqrt{1-\lambda^{2}}[f_{0}(x-yI)-f_{0}(x+yI)]}{2}$
$=G_{0}+KG_{1},\qquad G_{0},G_{1}\in\mathbb{B}\cap\mathbb{C}_{I}.$
Hence we have an orthogonal splitting and by a simple calculation
$|f(x+yJ)|^{2}=|G_{0}|^{2}+|G_{1}|^{2}=\left(\frac{1+\lambda}{2}\right)|f_{0}(x-yI)|^{2}+\left(\frac{1-\lambda}{2}\right)|f_{0}(x+yI)|^{2}.$
The last term is clearly bounded above by $\max\{|f_{0}(x-yI)|^{2},|f_{0}(x+yI)|^{2}\}$, which by our assumption is strictly less than $1$. This holds for all $J\in\mathbb{S}$ and therefore $f$ is a self map of the unit ball which is slice regular and solves the interpolation problem.
When $|Q_{n-1}^{n}|<1$, by [6, Thm. 4.1] $f_{0}$ varies in an infinite family of holomorphic functions, giving rise to infinite slice regular functions from $\mathbb{B}$ to $\mathbb{B}$ satisfying the requested interpolation. If $|Q_{n-1}^{n}|=1$, then $f_{0}$ is unique and given by a complex Blaschke product of degree $\kappa\in\{1,\ldots,n-1\}$:
$f_{0}(\zeta)=\left(\prod_{\ell=1}^{\kappa}\frac{\zeta-q_{\ell}}{1-\zeta\overline{q_{\ell}}}\right)\cdot v,\quad\zeta,q_{1},\ldots,q_{\kappa}\in\mathbb{B}\cap\mathbb{C}_{I},v\in\partial\mathbb{B}.$
Since $q_{1},\ldots,q_{\kappa}$ belong to the same slice $\mathbb{C}_{I}$, then the regular Blaschke product $B=\mathcal{M}_{q_{1}}\ast\ldots\ast\mathcal{M}_{q_{\kappa}}v$ is such that $B\mid_{\mathbb{C}_{I}}=f_{0}$: hence the unique extension of $f_{0}$ to a slice regular self-map of the unit ball $\mathbb{B}$ is exactly $B$ and we get that a solution of the interpolation problem is given by a regular Blaschke product of degree less or equal than $n-1$, which preserves the slice $\mathbb{C}_{I}$. Any solution preserving that slice must coincide with $B$ on the slice and hence everywhere.